### v1

# Methods and Open Problems in Differentiable Social Choice: Learning Mechanisms, Decisions, and Alignment

Abstract
Social choice is no longer a peripheral concern of political theory or economics—it has become a foundational component of modern machine learning systems. From auctions and resource allocation to federated learning, participatory governance, and the alignment of large language models, machine learning pipelines increasingly aggregate heterogeneous preferences, incentives, and judgments into collective decisions. In effect, many contemporary machine learning systems already implement social choice mechanisms, often implicitly and without explicit normative scrutiny.
Classical social choice and mechanism design offer deep axiomatic foundations for collective decision-making, but their traditional formulations are poorly matched to the scale, adaptivity, and distributional complexity of today’s socio-technical systems. In response, a rapidly growing literature has begun to reframe social choice, mechanism design, and alignment as differentiable learning problems, in which voting rules, allocation mechanisms, and incentive schemes are parameterized as neural or differentiable architectures, optimized from data, and evaluated empirically.
This Review synthesizes this emerging paradigm—what we term differentiable social choice—across auctions, voting rules, participatory budgeting, liquid democracy, decentralized aggregation, AI alignment, and inverse mechanism learning. We articulate a unifying perspective in which loss functions act as implicit aggregation rules, architectural inductive biases encode axioms such as anonymity or feasibility, and properties like robustness, proportionality, and incentive compatibility emerge as learned—but fragile—outcomes. We show how classical impossibility results and axiomatic trade-offs do not disappear under learning, but are relocated into objectives, constraints, data collection protocols, and optimization dynamics.
Beyond synthesizing methods, this Review is explicitly forward-looking. Across six domains, we identify 36 concrete open problems spanning incentive guarantees, robustness to distribution shift and strategic manipulation, certification and auditability of learned mechanisms, pluralistic preference aggregation, governance of alignment objectives, and the limits of inverse inference from behavior. These open problems delineate a new research agenda at the intersection of machine learning, economics, and democratic theory.
We argue that as machine learning systems increasingly mediate collective decisions at societal scale, the success of data-driven aggregation will depend not only on empirical performance, but on our ability to specify, verify, and govern the social choice principles that learning systems inevitably implement.
1 Introduction
The intersection of artificial intelligence and social choice theory is undergoing a structural transformation. For much of the twentieth century, collective decision-making—elections, public good provision, auctions, and committee decisions—was treated as an axiomatic enterprise. Mechanisms were derived as closed-form objects from normative desiderata such as strategyproofness, Pareto efficiency, anonymity, or revenue optimality. This classical program yielded foundational insights and impossibility results, including Arrow’s theorem, Gibbard–Satterthwaite, and Myerson’s optimal auction theory Arrow (2020); Gibbard (1973); Satterthwaite (1975); Myerson (1981). At the same time, these results expose intrinsic limitations of the axiomatic approach. Many desirable properties are mutually incompatible, guarantees are typically worst-case and agnostic to the statistical distribution of real-world data, and the resulting mechanisms often rely on strong rationality and tractability assumptions that are violated in practice Hart et al. (2013); Caiata et al. (2025).
Modern deployments have amplified these limitations. Spectrum auctions, online advertising markets, participatory budgeting, decentralized governance platforms, and preference-based training of large language models (LLMs) all operate in decision spaces that are high-dimensional, combinatorial, non-convex, and distribution-dependent. Yet, social choice is (both implicitly and explicitly) used in many of the most important contemporary machine learning research, e.g. in automated auction mechanism design for revenue optimzation Duan et al. (2023) and Large Language Models Duetting et al. (2024), in stake-holder portfolio balance and participatary resource allocation Fairstein et al. (2024), in federated learning where model updates are aggregated based on a certain mechanism Kesari & Kesari (2025), and, most celebrated recently, in preference-based alignment (RLHF) of LLMs from human annotators who have inherently heterogeneous preference models Siththaranjan et al. (2024); Swamy et al. (2024). We summarize these applications of social choice in machine learning in Table 1. In such settings, analytic derivations with universal guarantees can be simultaneously too pessimistic—ruling out mechanisms that perform well on realistic data—and too optimistic—failing catastrophically under distribution shift or strategic adaptation Hart et al. (2013); Caiata et al. (2025). Nevertheless, many contemporary applied systems are already learned: allocation policies, ranking rules, and reward models are trained by gradient descent, whether or not designers acknowledge their implicit social-choice content.
This tension has catalyzed the emergence of Differentiable Social Choice. The central premise is that mechanisms need not be static artifacts of logic; instead, they can be parameterized, differentiable functions optimized from data, subject to explicit structural and normative constraints. Drawing on advances in differentiable economics Dütting et al. (2017; 2019); Anil & Bao (2021), neural social choice Matone et al. (2024); Hornischer & Terzopoulou (2024), and learning-based mechanism design Fairstein et al. (2024); Kesari & Kesari (2025), this paradigm treats voting rules, auctions, and allocation mechanisms as trainable objects. Classical axioms are no longer enforced solely by analytical proofs; they reappear as inductive biases, architectural symmetries, loss terms, and audit criteria.
Concrete applications already illustrate both the promise and the stakes of this shift. In auction design, neural mechanisms have been trained to approximate revenue-optimal auctions in multi-item settings where analytic solutions are unknown, encoding incentive compatibility as regret penalties or architectural constraints Dütting et al. (2019); Ivanov et al. (2022); Hertrich et al. (2023). In participatory budgeting, learned aggregation rules operate over knapsack-feasible portfolios, optimizing proportionality and efficiency objectives at scales where classical solvers struggle, while benchmarking against principled rules such as the Method of Equal Shares (MES) Peters et al. (2021); Fairstein et al. (2024). In liquid democracy, delegation graphs are implemented as differentiable aggregation architectures, directly connecting social choice to federated learning and edge aggregation under trust constraints Kesari & Kesari (2025); Alouf-Heffetz et al. (2025). In AI alignment, preference-learning objectives used to train reward models and policies have been shown to implicitly instantiate specific voting rules—most notably Borda count—even when designers intend to remain agnostic about normative aggregation Siththaranjan et al. (2024); Dai & Fleisig (2024). More recently, inverse mechanism learning has turned these tools toward governance and accountability, aiming to reconstruct opaque incentive structures from observed strategic behavior An & Du (2026); Ahmadi et al. (2023).
These developments mark a deep conceptual shift. We structurally illustrate this conceptual shift in Figure 1. Differentiable social choice does not abandon classical theory; it repurposes it. Social choice axioms become a specification language for loss functions and architectures, and an audit vocabulary for interpreting learned systems. Anonymity and neutrality translate into permutation-equivariant neural networks; feasibility and budget constraints become differentiable solver layers; incentive compatibility is relaxed into approximate regret bounds that can be empirically evaluated and stress-tested. At the same time, the differentiable turn exposes new risks and novel research gaps: soft constraints may fail silently, expressive models may hide violations, and optimization objectives may encode unintended normative commitments.
Figure 1: A structural shift from axiomatic social choice (closed-form mechanisms derived from normative axioms) to differentiable social choice (parameterized mechanisms optimized from data), where axioms reappear as architectural biases, loss terms, and audit criteria.
Table 1: Where social choice appears in modern ML systems. Many modern ML pipelines combine signals from multiple agents or sources in ways that are mathematically analogous to social choice or mechanism design principles. The “implicit aggregation rule” column points to how such aggregation arises in practice, and the final column gives representative references supporting these connections.
ML system Agents / sources Aggregated signal Implicit aggregation rule Representative reference
Learned auction design Strategic bidders Bids, valuations Optimization of revenue under incentive constraints creates an implicit choice over allocation and payment functions Automated auctions using neural networks and mechanism learning Duan et al. (2023); Dütting et al. (2021)
Neural voting / ranking models Voters or users Rankings, approvals, clicks Learning to predict or rank outcomes aggregates individual preferences similar to positional and Condorcet rules Chen et al. (2023); Boutilier et al. (2023); Suryanarayana et al. (2022)
Data-driven PB aggregation Citizens / stakeholders Approvals or utility vectors Machine learned selection of portfolios balances utility and representation, analogous to PB rules Learning aggregation rules in participatory budgeting (data-driven PB) Fairstein et al. (2024)
Federated learning / decentralized aggregation Distributed clients Local model updates Weighted or consensus-based aggregation corresponds to a voting/collective influence mechanism Liquid democracy as hybrid voting model (proxy delegation) Kesari & Kesari (2025)
Preference-based alignment (RLHF) Human annotators Pairwise preference labels Aggregating pairwise comparisons resembles social choice rules such as Borda/Braŋdley-Terry mechanisms Mapping social choice theory to RLHF and preference aggregation Siththaranjan et al. (2024); Dai & Fleisig (2024); Conitzer et al. (2024); Swamy et al. (2024)
Platform governance / incentive forensics Strategic users Behavioral adaptation traces Inferring effective rules from observed strategic behavior is analogous to inverse mechanism design IC and mechanism design for sociotechnical systems in AI alignment Zhang et al. (2024); An & Du (2026)
This Review surveys this rapidly evolving landscape with a unifying perspective. We organize the literature into six interconnected domains: (1) Differentiable Economics, which studies learning-based approximations to optimal auctions and contracts; (2) Neural Social Choice, which synthesizes, analyzes, and audits voting rules using deep learning; (3) Participatory Budgeting, where AI systems support collective resource allocation under combinatorial constraints; (4) Liquid Democracy, linking delegation and decentralization to graph-based learning; (5) AI Alignment as Social Choice, which interprets preference learning and reinforcement learning from human feedback (RLHF) as implicit voting; and (6) Inverse Mechanism Learning, which seeks to infer and audit incentive structures from behavior.
Our goal is not merely to catalog methods, but to articulate a coherent narrative: the transition from axiomatic artifacts to learned institutions. As collective decision-making becomes embedded in machine learning pipelines, the design of loss functions, architectures, and evaluation protocols becomes inseparable from democratic theory. Differentiable social choice thus emerges not only as a technical research program, but as a foundational framework for the governance, auditability, and legitimacy of algorithmic decision-making systems.
2 Differentiable Economics: From Optimal Auctions to Distributed Edge Markets
Mechanism design is often described as “inverse game theory”: rather than predicting equilibria of a given game, the designer chooses the game form so that equilibrium behavior induces a desirable outcome. In the classical theory, this program is driven by structural characterizations and closed-form solutions—most prominently, Myerson’s optimal auction for a single item with independent private values Myerson (1981). Yet, even in stylized multi-item settings, optimal mechanisms can require high-dimensional (and potentially exponentially complex) menus, intricate randomization, and discontinuous allocation rules Hart et al. (2013); Hartline & Roughgarden (2009). This gap between elegant theory and messy, high-dimensional practice has made automated mechanism design an enduring ambition in economics and computer science Conitzer & Sandholm (2002); Sandholm (2003).
Differentiable economics reframes this ambition as a learning problem: parameterize the allocation and payment rules as differentiable function approximators (typically neural networks), optimize expected revenue or welfare over samples from a valuation distribution, and enforce incentive constraints via differentiable surrogates Dütting et al. (2017; 2019). The conceptual shift is not merely computational convenience. It changes the object of study from axioms $\rightarrow$ closed-form mechanisms, to objectives + constraints $\rightarrow$ learned mechanisms whose structure is discovered (and sometimes audited) empirically. This section reviews the methodological core, the rapidly evolving architectural choices (from multilayer perceptrons to attention and graph-based inductive biases), the emerging theory for optimization and verification, and the expansion from textbook auctions to resource-constrained, distributed edge markets.
2.1 From analytical optimal auctions to learned mechanisms
Classical foundations and why multi-dimensionality is hard.
In single-parameter environments (e.g., one item, or matroid constraints with single-dimensional types), incentive compatibility admits strong structural characterizations; for example, monotone allocations and threshold payments yield dominant-strategy incentive compatibility (DSIC) Myerson (1981). In contrast, in multi-parameter environments (multiple heterogeneous items, bundles, or quality levels), the feasible set of DSIC mechanisms is governed by cyclic monotonicity of the allocation rule, and payments correspond to (sub)gradients of a convex potential Rochet (1987). This geometry is powerful but rarely yields closed forms outside special cases. Theoretical work shows that optimal multi-item mechanisms can be highly complex (sometimes requiring large menus or delicate randomization), making the search space difficult to reason about analytically Hart et al. (2013).
Neural parameterizations: allocation and payment networks.
The differentiable approach implements a mechanism as two coupled maps: an allocation network $x_{\theta}(\cdot)$ producing (possibly fractional or randomized) allocations, and a payment network $p_{\theta}(\cdot)$ producing transfers. The earliest influential line, often associated with the “deep automated mechanism design” paradigm, trains these networks end-to-end to maximize expected revenue under incentive constraints Dütting et al. (2017; 2019). In practice, the allocation network is often designed to satisfy feasibility by construction (e.g., softmax-based assignment for unit-demand, differentiable sorting/ranking for position auctions, or projection layers for budgets), while payments are produced either by a separate network or by integrating along value paths when exploiting Rochet-style structure Dütting et al. (2017; 2019); Rochet (1987).
Recovering known solutions and discovering new ones.
A key empirical validation is that learned mechanisms can recover canonical solutions where theory is complete (e.g., Myerson-like reserve pricing in single-item auctions) and outperform strong baselines in settings where the optimal mechanism is unknown Dütting et al. (2017; 2019). This is not simply curve-fitting: the learned mapping must simultaneously satisfy feasibility, approximate DSIC/BIC, and optimize revenue, which tends to expose where classical analytical templates (e.g., simple itemwise pricing) leave money on the table in correlated or multi-item settings Dütting et al. (2019); Huo et al. (2025).
2.2 Differentiable incentive constraints: regret, cyclic monotonicity, and certification
Regret minimization as a differentiable surrogate.
The dominant technical obstacle is that incentive constraints are universal quantifications over deviations. For DSIC, truthful reporting must maximize utility for every agent and every profile of others’ bids. Regret-based training replaces this with a differentiable penalty:
$\mathrm{Regret}_{i}(\theta;v)\;=\;\max_{\hat{v}_{i}}\;u_{i}\!\big(\hat{v}_{i},v_{-i};\theta\big)\;-\;u_{i}\!\big(v_{i},v_{-i};\theta\big),$ (1)
and optimizes expected revenue subject to low expected regret, often via Lagrangian or augmented Lagrangian schemes Dütting et al. (2017). This family of methods is frequently associated with RegretNet-style pipelines, in which an inner maximization (adversarial misreport search) is approximated by gradient ascent on $\hat{v}_{i}$, while an outer minimization updates mechanism parameters Dütting et al. (2017).
A subtlety, highlighted by subsequent work, is that regret estimation itself can be unreliable: the inner maximization is non-convex and can under-estimate true regret, producing mechanisms that appear incentive compatible during training but violate constraints under stronger attacks You et al. (2026). This has motivated better adversarial optimizers, validation protocols tailored to regret-based mechanisms, and objective formulations that expose a more interpretable trade-off between revenue and acceptable IC violation (e.g., regret budgets) Ivanov et al. (2022); You et al. (2026).
RochetNet and structure from cyclic monotonicity.
An alternative to pure regret penalization is to enforce DSIC via structural conditions derived from convex analysis. Rochet’s theorem provides a necessary and sufficient condition for implementability in quasilinear multi-parameter domains: an allocation is DSIC implementable iff it is cyclically monotone, and payments can be derived from a convex potential Rochet (1987). RochetNet-style constructions exploit this by parameterizing a convex function and defining allocations/payments through its gradients, achieving architectural incentive properties rather than purely penalty-based ones Dütting et al. (2019); Hertrich et al. (2023). In practice, such approaches trade expressivity and optimization simplicity against stronger built-in guarantees.
Certification and verification: beyond empirical audits.
As differentiable mechanisms move toward deployment, the field has begun to explore certification: given a trained network, can we prove it is DSIC (or approximately DSIC within a bound) over a domain? “Certifying strategyproof auction networks” develops tooling for post-training verification in restricted settings, reflecting a broader push to reconcile black-box learning with the normatively central requirement of incentive guarantees Curry et al. (2020). Verification remains challenging because IC is a global property over a continuous type space and can be brittle under distribution shift or stronger deviation models than those used in training You et al. (2026).
Table 2 summarizes the main approaches to enforcing incentive compatibility in differentiable mechanism design, highlighting how they differ in enforcement mechanism, strength of guarantees, computational cost, and characteristic failure modes.
Table 2: Incentive enforcement strategies in differentiable mechanism design. Methods differ in where incentive constraints enter (training-time penalties vs. architectural implementability vs. post-hoc certification), the strength of guarantees (empirical expected regret vs. structural/certified bounds), and characteristic failure modes (notably regret underestimation and distribution-shift brittleness).
Approach IC target Enforcement mechanism Optimization structure Guarantee type Compute profile Expressivity Typical failure modes / best use
Regret penalties Dütting et al. (2017; 2019) Approx. DSIC / BIC Add regret term $\mathbb{E}[\max_{\hat{v}_{i}}u_{i}(\hat{v}_{i})-u_{i}(v_{i})]$; inner-loop deviation search Bilevel min–max; Lagrangian / ALM Empirical / in-distribution (expected regret on samples) High (inner maximization per batch) High (generic NN mechanisms) Failure: regret underestimation when the inner maximizer is weak; IC can break under stronger attacks or shift You et al. (2026). Best: flexible baseline for high-dimensional auctions.
Stronger deviation solvers Ivanov et al. (2022); You et al. (2026) Approx. DSIC / BIC Improve inner maximization (multi-start, longer ascent, better attacks); sometimes regret budgets Still bilevel; often tighter outer objectives Still empirical (but more reliable stress-tests) Very high (more attack steps / restarts) High Failure: remains attacker-class dependent; can overfit to chosen deviation model. Best: when IC violations are costly and compute is available.
Architectural IC (cyclic monotonicity / convex potential) Rochet (1987); Hertrich et al. (2023) DSIC implementability (domain-dependent) Parameterize convex potential; derive allocation/payment via gradients to satisfy implementability conditions Single-level training (no explicit inner max) Structural (architecture-level, within modeled domain) Moderate Medium (restricted class) Failure: mismatch to domain assumptions; limited mechanism family; feasibility layers can still be brittle. Best: when built-in IC is required and structure is acceptable.
Structured mechanism templates Li et al. (2024) Template-dependent (often approximate IC) Constrain to interpretable families (e.g., rank-score rules, parametric payments) and learn parameters Single-level ERM / decision-focused Partial / template-inherited Low–moderate Low–medium Failure: may leave revenue/welfare on the table; hidden violations if template assumptions fail. Best: deployments needing interpretability, regulation fit, or low latency.
Post-hoc certification / verification Curry et al. (2020) DSIC or bounded approx. DSIC Verify trained network over a domain (e.g., partitions / discretization / bounds) Train-then-certify pipeline Certified bound on checked domain Potentially high (scales with dimension/coverage) Any (but certifiable only in restricted settings) Failure: scalability and coverage gaps; guarantees may not extend beyond tested region. Best: high-stakes auditing and compliance.
Robustness-aware training (distributional / ambiguity) Ball & Kattwinkel (2024) IC + performance under shift Optimize worst-case or neighborhood objectives; ambiguity sets / stability refinements Min–max or robust ERM variants Model-dependent robustness guarantees High–variable Medium–high Failure: misspecified ambiguity sets can induce performance cliffs; robustness can trade off sharply with revenue. Best: settings with known drift/uncertainty in bidder distributions.
2.3 Architectures and inductive biases: symmetry, sets, and attention
Why inductive bias matters in economic learning.
Mechanisms are typically anonymous (invariant to permuting bidder identities) and often neutral with respect to symmetric items. Generic multilayer perceptrons must learn these symmetries from data, wasting capacity and risking spurious identity leakage. This motivates architectures that enforce permutation equivariance/invariance by construction Rahme et al. (2021).
Permutation-equivariant RegretNet variants.
A prominent line introduces permutation-equivariant layers for auction networks, showing both sample-efficiency and revenue gains by baking symmetry into the model class Rahme et al. (2021). “Optimal-er Auctions through Attention” advances this theme by replacing MLP blocks with attention modules (RegretFormer) and proposing a regret-budgeted loss that makes the revenue–IC trade-off more controllable Ivanov et al. (2022). Attention is not only a modern deep learning convenience; it offers a principled way to represent auctions as set-to-set mappings where interactions among bidders and items must be modeled flexibly while respecting exchangeability.
Transformers for combinatorial and joint auctions.
Recent work pushes beyond bidder symmetry to the deeper challenge of combinatorial structure. In joint auctions (e.g., simultaneous allocation of multiple related opportunities), both cross-bidder and cross-item dependencies can be high-order and context dependent. Transformer-based designs such as JTransNet (and related CAFormer/CANet-style variants) use self-attention to model these interactions, naturally handling variable-sized sets and enabling anonymous deterministic allocation in joint auction settings Zhang et al. (2025). Empirically, these approaches often dominate older baselines in revenue while improving scalability across varying numbers of bidders/items, at the cost of increased training complexity and the need for careful constraint handling Zhang et al. (2025).
Domain-guided architectures in ad auctions and industrial marketplaces.
A complementary industrial thread constrains the mechanism class to improve interpretability and deployment fit. For example, rank-score allocation rules and second-price-style payments can be parameterized and learned end-to-end to optimize platform metrics in e-commerce advertising, producing mechanisms that remain close to Generalized Second Price templates while benefiting from data-driven tuning Li et al. (2024). This “structured learning” view suggests a continuum between fully flexible neural mechanisms and tightly constrained parametric families—with the right point on the continuum determined by deployment needs (transparency, latency, regulation) as much as by revenue.
2.4 Optimization landscapes and theory: generalization, connectivity, and expressivity
Generalization bounds and the role of sampling.
Because objectives are expectations over valuation distributions, differentiable economics naturally invites statistical learning theory: how many samples are needed for near-optimal revenue, and how does approximation error scale with network capacity and constraint enforcement? Early deep auction design work provides formal generalization bounds under regret-based formulations, clarifying when learned mechanisms should transfer from training draws to unseen valuations Dütting et al. (2017; 2019). These results are only partial—they typically bound performance under the assumed distribution and do not, by themselves, guarantee robustness to misspecification.
Mode connectivity and benign non-convexity.
A persistent concern is that training is non-convex and could converge to poor local optima. “Mode Connectivity in Auction Design” provides a theoretically grounded counterpoint: for RochetNet-like architectures (and generalized affine maximizer auctions), local minima are connected by simple low-loss paths in parameter space, implying that the effective landscape is surprisingly benign Hertrich et al. (2023). This connects differentiable mechanism design to a broader deep learning observation: many seemingly hard non-convex objectives exhibit connectivity among good solutions, which helps explain why stochastic gradient methods can reliably find high-quality mechanisms despite limited theoretical guarantees.
Expressivity meets economic complexity.
From an economic perspective, benign optimization does not eliminate the deeper question: what class of mechanisms can a given architecture represent? Classical work shows that optimal mechanisms over general type spaces may be discontinuous, menu-complex, or rely on subtle randomization Hart et al. (2013). Contemporary theory on general type spaces further emphasizes that even defining the right structural template can be delicate when types are disconnected or non-convex Prasad et al. (2025). Neural networks are attractive precisely because they can approximate complex mappings, but the tension remains between expressivity (capturing the true optimum) and verifiability (certifying IC, IR, fairness, and feasibility).
2.5 Robustness and strategic adaptivity: from distribution shift to adversarial learning
Distribution shift as a first-class economic concern.
A learned mechanism is only as reliable as the distributional assumptions implicit in its training data. If bidder populations, correlations, or value supports change, revenue and incentive properties can degrade sharply. In economics, this motivates robust mechanism design: optimize performance under ambiguity about the prior. “Robust Robustness” refines this perspective by observing that a mechanism designed for a worst-case ambiguity set can itself be fragile to misspecification of that set; it proposes a robustness notion that requires payoff guarantees to be stable in neighborhoods around the ambiguity set (in the weak topology), avoiding “performance cliffs” Ball & Kattwinkel (2024). This refinement resonates strongly with modern machine learning concerns about out-of-distribution generalization and adversarially induced failures.
Strategic manipulation of the learning process.
Beyond passive shift lies an active threat: agents may strategically manipulate not only their bids within a fixed mechanism, but also the seller’s learning of the mechanism over time. “Adversarial Learning in Revenue-Maximizing Auctions” studies repeated auctions where sellers adapt mechanisms based on past bids; bidders can learn bidding strategies that exploit this adaptivity even without full knowledge of the seller’s update rule, improving bidder utility and potentially reducing seller revenue Nedelec et al. (2019). This perspective reframes differentiable economics as a learning-in-games system: the seller runs gradient descent on mechanism parameters while bidders run their own learning dynamics, producing a coupled non-stationary environment. It also suggests an emerging security agenda: mechanism learning algorithms should be robust to strategic data poisoning by bidders, analogous to adversarial training in machine learning Nedelec et al. (2019); You et al. (2026).
Correlated values and richer priors.
Much of the early neural auction literature uses independent priors for tractability, but real markets commonly exhibit correlation (shared signals, common-value components, contextual effects). Learning-based approaches can incorporate such structure directly by sampling from correlated distributions and optimizing end-to-end, though incentive and identification issues can become more subtle Huo et al. (2025). This intersects with a broader trend in mechanism design: moving from analytically convenient priors to data-driven, high-dimensional priors inferred from platform logs.
2.6 From auctions to edge markets: partial fulfillment and resource-constrained mechanism design
Why edge markets change the mechanism design problem.
Edge computing markets allocate heterogeneous resources (CPU, GPU, memory, bandwidth, energy) across distributed service providers and application service providers. Unlike canonical item auctions, allocations can be divisible and partially fulfilled: an agent may receive 60% of requested compute, with utility and feasibility governed by engineering constraints, latency, and network topology. Classical mechanisms such as VCG Clarke (1971); Vickrey (1961); Groves (1973) may be computationally burdensome or misaligned with partial fulfillment and operational constraints.
Neural mechanisms for partial fulfillment.
“Neural Auction Mechanisms for Partial Fulfillment in Edge Computing” proposes architectures explicitly tailored to this domain, decoupling allocation decisions into probability and quantity components (PDNet) and using exchangeable attention mechanisms (EANet) to preserve bidder anonymity while scaling across variable numbers of agents Zhang et al. (2026). The learned mechanisms are trained with revenue objectives and regret-based incentive penalties, while embedding feasibility constraints relevant to edge capacity and service provisioning. A notable contribution is the modularity of the differentiable approach: by altering output parameterizations and constraint projections, the same learning framework can address allocation regimes that are awkward for traditional all-or-nothing auction templates Zhang et al. (2026).
Distributed implementation and the path to “edge-native” mechanism design.
Edge settings also raise systems questions often absent in classical auction theory: decentralized information, limited communication, privacy constraints, and real-time decision requirements. These pressures motivate mechanisms that are lightweight to evaluate, stable under asynchronous updates, and robust to missing or delayed signals. While current work largely trains centralized networks and deploys them as decision policies, the architectural trend toward permutation-equivariant and set-based models is compatible with more distributed realizations (e.g., message passing or hierarchical attention), hinting at a convergence between differentiable mechanism design and distributed learning.
2.7 Open Problems
OP1: Hard incentive guarantees in expressive networks.
How can we move from soft regret penalties and empirical audits to hard, architecture-level DSIC (or certified approximate DSIC) guarantees for high-capacity models without collapsing expressivity?
OP2: Reliable regret estimation under strong deviations.
Can we design inner-loop deviation solvers and validation protocols that provably upper bound regret (rather than under-estimate it) in realistic continuous domains, especially for transformer-based mechanisms?
OP3: Robustness to distribution shift and ambiguity misspecification.
What is the right unification of distributionally robust mechanism design (including refinements such as robust robustness) with modern out-of-distribution learning so that learned mechanisms degrade gracefully under shifting bidder populations?
OP4: Mechanism learning under strategic data poisoning.
How do we design seller-side learning dynamics that are provably robust when bidders strategically manipulate bids to influence the future mechanism, rather than only optimizing within a fixed mechanism?
OP5: Edge-native decentralization and communication constraints.
Can differentiable mechanisms be trained and deployed in a decentralized manner that respects edge communication limits and privacy constraints while maintaining feasibility and incentive properties?
OP6: Normative constraints beyond revenue.
How should we represent and enforce normative desiderata—fairness, diversity, market access, or sustainability—in differentiable objectives so that these constraints are not merely post-hoc metrics but are satisfied reliably at deployment time?
3 Neural Social Choice and Voting Rules
Social choice theory studies how to aggregate individual preferences into a collective decision. In its classical form, the subject is defined by axioms (normative desiderata) and impossibility results: Arrow’s theorem shows that no rank-based social welfare function can simultaneously satisfy a natural set of fairness and rationality axioms once there are at least three alternatives Arrow (2020); Sen (1986). In the same spirit, Gibbard–Satterthwaite establishes that every onto, non-dictatorial deterministic voting rule with at least three alternatives is manipulable Gibbard (1973); Satterthwaite (1975). These foundational barriers gave rise to a taxonomy of voting rules—plurality, Borda, Condorcet extensions, approval, STV, and many others—each embodying an explicit compromise between incompatible ideals Black (1958); Fishburn (1974); Tideman (2006).
Neural social choice (and, more broadly, differentiable social choice) asks a different design question: rather than selecting among a small menu of named rules, can we learn an aggregation rule from data, while steering it toward an intended axiomatic compromise? This move is motivated by three pressures. First, many attractive rules are computationally hard to compute exactly (e.g., Dodgson, Young, Kemeny–Young), making them difficult to deploy at scale Bartholdi et al. (1989); Conitzer et al. (2007). Second, worst-case axiomatic analysis can be overly pessimistic: rule failures may be rare or common depending on the distributional structure of preferences (e.g., single-peakedness, Mallows noise, or real-world correlations) Marden (1995); Procaccia (2010); Elkind et al. (2016). Third, modern collective decisions increasingly happen in algorithmic settings (recommenders, ranking, RLHF, and platform governance) where the “electorate” and its preferences are observed as data streams rather than as a one-off election Gleich & Lim (2011); Mattei & Walsh (2013).
This chapter reviews the emerging literature around (i) representation and architectural inductive bias for preference profiles, (ii) mimicry of classical rules versus synthesis of new rules, (iii) axiomatic auditing and the generalization gap between outcome imitation and principled behavior, (iv) distribution-specific evaluation of rule violations in practice, (v) robustness and adversarial training for manipulation resilience, and (vi) “loss-as-rule” frameworks that make the aggregation principle explicit and differentiable.
3.1 Axioms, impossibilities, and the learning objective
From binary axiom satisfaction to differentiable proxies.
Classical voting axioms are typically discrete properties of a function $f$ mapping a preference profile to an outcome (winner, ranking, or committee). Core examples include anonymity (voter symmetry), neutrality (candidate symmetry), Pareto efficiency, monotonicity, Condorcet consistency, and (for multi-winner rules) proportionality and representation guarantees Moulin (1988a); Young (1995); Brandt et al. (2016). In neural settings, these axioms are often evaluated empirically as violation rates over sampled profiles, or encoded as differentiable penalties that approximate discontinuous desiderata Hornischer & Terzopoulou (2024); An et al. (2026).
A central conceptual shift is that the object being designed is no longer a closed-form rule with a known proof portfolio, but a function class (e.g., a permutation-invariant network) optimized against a loss. This reframes axioms as (i) constraints to be satisfied (hard or soft), (ii) terms in a multi-objective optimization, or (iii) evaluation metrics that can sharply diverge from the training objective. The resulting design space is closer to ML model selection than to classical axiomatic deduction, raising questions about transparency, validation, and normative governance.
Manipulation as a learning-and-security problem.
Gibbard–Satterthwaite is not nullified by deep learning: any sufficiently expressive deterministic rule that is non-dictatorial and onto remains manipulable in principle Gibbard (1973); Satterthwaite (1975). What changes is the empirical and operational framing. In learned rules, we can: (i) measure attack surfaces and attack success rates under bounded coalitions or bounded vote changes, (ii) train with adversarial examples to reduce these rates, and (iii) trade off manipulation resistance against other desiderata (e.g., welfare) via explicit objectives Li et al. (2026); Procaccia (2010).
3.2 Representing preference profiles: symmetry, sets, and canonical embeddings
Why representation is the first mechanism.
A preference profile for $n$ voters and $m$ candidates can be encoded as (i) a matrix of ranks, (ii) pairwise comparisons, (iii) positional score vectors, or (iv) a multiset of permutations. In learning, the encoding implicitly imposes structure: a naive one-hot tensor with fixed voter indices risks breaking anonymity; similarly, candidate indexing can leak spurious biases if neutrality is not respected. This motivates permutation-invariant/equivariant architectures as a form of architectural axiom enforcement Zaheer et al. (2017); Lee et al. (2019).
Permutation-invariant neural social choice.
Early work already observed a surprising phenomenon: standard neural classifiers trained to predict winners often behave like simple positional rules (notably Borda), even when trained on other targets, suggesting strong inductive biases induced by common encodings and optimization dynamics Burka et al. (2016); Armstrong (2021). “Learning to Elect” systematizes this by casting voting rules as set-input functions and comparing DeepSets, Set Transformers, and graph networks for (a) mimicking classical rules and (b) discovering welfare-maximizing rules Anil & Bao (2021). The key takeaway is that set-invariant models offer both theoretical alignment with anonymity/neutrality and practical gains in sample efficiency and generalization.
Canonical embeddings from social choice theory (DeepVoting).
DeepVoting pushes representation further by importing canonical embeddings from social choice into the learning pipeline Matone et al. (2024). Instead of learning directly from raw rank matrices, DeepVoting uses embeddings derived from positional scoring vectors and related structures, and recasts deterministic rules as probabilistic social choice functions that output distributions over candidates. This probabilistic relaxation is not merely an ML trick: it smooths discontinuities, supports gradient-based fine-tuning for normative properties, and naturally aligns with settings where collective decisions are inherently stochastic (e.g., randomized tie-breaking, sampling-based committees, or probabilistic recommendation). A central empirical claim is that tailored embeddings reduce model size requirements and improve scaling to larger electorates relative to prior black-box encodings Matone et al. (2024).
3.3 Mimicry versus synthesis: what does it mean to “learn a voting rule”?
Mimicking hard-to-compute rules.
One pragmatic motivation for neural social choice is computational: approximate a rule that is normatively appealing but costly to compute (e.g., Kemeny–Young, Dodgson, Young) by a fast forward pass Bartholdi et al. (1989); Conitzer et al. (2007). “Learning to Elect” demonstrates that modern set-invariant networks can mimic both positional (plurality, Borda) and more complex rules at high accuracy, with generalization depending critically on distribution shift and profile size changes Anil & Bao (2021). DeepVoting similarly targets efficient learning of complex rules through structured embeddings Matone et al. (2024). This line is valuable as an approximation infrastructure—but it also exposes a conceptual pitfall: matching outputs on typical training distributions does not imply the learned function satisfies the axioms that define the target rule.
Synthesis: learning rules that optimize welfare or axioms directly.
A second, more radical agenda is to treat the rule itself as a decision policy optimized for a population and a criterion. “Learning to Elect” trains networks to maximize utilitarian or egalitarian welfare under assumed voter utility models, and reports learned rules that outperform classical ones on the target welfare measure, matching known optima in special cases Anil & Bao (2021). DeepVoting extends this idea through axiomatic fine-tuning: a network pretrained to mimic a standard rule can be updated (via gradient descent) to reduce pathologies such as a probabilistic version of the no-show paradox, effectively “patching” a rule’s weaknesses while keeping much of its behavior intact Matone et al. (2024); Moulin (1988b). This illustrates a distinctive promise of differentiable design: rather than choosing between Borda and Condorcet families, one can navigate a continuous space of hybrid rules between them.
3.4 Axiomatic auditing and the “mimicry–reasoning” gap
Axiomatic Deep Voting (Learning How to Vote with Principles).
Hornischer and Terzopoulou provide a critical “reality check” by developing an axiomatic audit framework for neural voting rules Hornischer & Terzopoulou (2024). They report a disconnect between (i) high predictive accuracy in mimicking a target rule’s outcomes and (ii) low satisfaction of that rule’s defining axioms under distribution shift or out-of-distribution tests. Notably, they argue that training on axiom-specific data does not automatically repair axiom satisfaction. Instead, they show that directly optimizing axiom satisfaction (through axiom-derived loss terms) can synthesize novel neural rules that occupy new regions of the axiom trade-off landscape and can substantially differ from any named classical rule Hornischer & Terzopoulou (2024).
Why the gap appears: discontinuity, rare constraints, and compositionality.
Several mechanisms plausibly contribute to the mimicry–reasoning gap. Many voting rules are piecewise-constant with complex decision boundaries; small perturbations near boundaries can flip winners, making generalization fragile in high-dimensional profile spaces. Moreover, axioms often constrain behavior on rare configurations (e.g., specific paradoxical profiles), so finite-sample mimicry may never “see” the regions where axioms bite. Finally, axioms are global and compositional (they quantify over counterfactual profiles), while supervised training is pointwise. The audit-first perspective thus becomes essential: learned democracies cannot be assumed to inherit normative properties from their labels Hornischer & Terzopoulou (2024).
Interpretability and explanation of learned rules.
A complementary response is to explain learned rules in human-legible terms. “Learning to Explain Voting Rules” learns interpretable surrogates (e.g., trees) that approximate voting behavior and can expose which profile features drive outcomes, supporting governance and debugging Kang et al. (2023). This connects neural social choice to the broader ML agenda of post-hoc interpretability, but with unusually crisp normative stakes: explanations are not only for trust but for verifying that the rule embodies the intended democratic compromise.
3.5 What voting rules actually do in practice: distribution-specific social choice
From worst-case theorems to empirical violation rates.
Much of social choice theory is worst-case: an axiom is either satisfied for all profiles or not. Caiata, Armstrong, and Larson argue that this binary view can be misleading for deployed systems that operate under structured preference distributions Caiata et al. (2025). They propose a data-driven evaluation framework for multi-winner voting rules that estimates how frequently axioms are violated across synthetic and real-world distributions. Their analysis suggests that classical rules with clean theoretical stories may violate desirable properties frequently on realistic distributions, while learned neural rules can be trained to reduce these violation rates for a target population Caiata et al. (2025).
Implications: personalization, legitimacy, and “bespoke” democracy.
The distribution-specific view suggests a provocative possibility: instead of universal rules, we may design electorate-adapted rules that are “more democratic” for a given society (in the sense of fewer violations of the society’s chosen axioms under its empirical preference structure). This resonates with empirical social choice and preference restriction models (e.g., single-peaked, single-crossing, Mallows) that already show how structure changes both feasibility and guarantees Marden (1995); Elkind et al. (2016). However, it also raises legitimacy questions: if rules are trained on historical data, which populations are represented, how robust is the rule to demographic change, and who gets to specify the axiom weights?
3.6 Robustness and manipulation resilience: adversarial learning for elections
Adversarial GNNs for resilient elections.
Li, Shah, and Giusti explicitly connect learned voting to AI security by proposing adversarially trained graph neural networks for elections Li et al. (2026). They represent an election as a bipartite graph (voters–candidates) and learn the aggregation function with a GNN, which naturally encodes anonymity and neutrality via permutation invariance over nodes. They introduce a min–max training regime in which an attacker perturbs votes (e.g., via strategic manipulation or noise injection) to flip outcomes, and the model is trained to maintain target properties such as welfare and robustness under these attacks Li et al. (2026).
Why graph structure is a natural inductive bias.
Graph representations unify several desiderata: (i) exchangeability (node permutation), (ii) variable-size generalization to different numbers of voters/candidates, and (iii) a clear interface for modeling coalitions or local perturbations as graph edits. This architecture choice aligns with earlier set-invariant approaches Zaheer et al. (2017); Lee et al. (2019); Anil & Bao (2021) but emphasizes attack modeling and robustness objectives. The broader research question is whether robustness metrics in elections should resemble adversarial robustness in classifiers, or whether domain-specific constraints (e.g., bounded coalition size, bounded Kendall distance) provide a more normatively meaningful threat model Procaccia (2010).
3.7 Differential Voting: loss functions as voting rules
Loss geometry encodes normative aggregation.
A striking bridge between neural social choice and ML practice is the observation that many learning pipelines implicitly implement voting rules. In preference-based learning (including RLHF reward modeling), the standard Bradley–Terry–Luce (BTL) log-likelihood corresponds to a particular aggregation principle closely related to Borda-like scoring An et al. (2026). Differential Voting makes this connection explicit by constructing differentiable, instance-wise losses whose population optima correspond to distinct classical rules, and by analyzing how loss geometry (margin sensitivity, boundary concentration) translates into axiomatic behavior An et al. (2026).
Differentiable surrogates for Condorcet-style aggregation.
Differential Voting proposes smooth surrogates for discontinuous criteria such as Copeland (pairwise majority wins) and Kemeny (minimizing pairwise disagreement), and establishes consistency and limiting behavior as smoothing vanishes An et al. (2026). Conceptually, this turns “choose a voting rule” into “choose a loss” and enables mix-and-match democracy: weighted sums of losses can interpolate between ideals (e.g., Condorcet consistency vs. smooth optimization stability). This approach also clarifies a central methodological tension in neural mechanism design: objectives that are normatively attractive often induce non-smooth or combinatorial optimization, while smooth objectives may bake in unintended normative biases.
3.8 Open Problems
OP7: Certification of axioms for learned voting rules.
How can we provide post-training certificates (or architecture-level guarantees) that a learned rule satisfies anonymity, neutrality, monotonicity, or Condorcet consistency beyond finite-sample testing?
OP8: Generalization under electorate shift.
What training objectives and validation protocols ensure that axiomatic behavior persists when the preference distribution changes (new candidates, new voter blocs, or different cultural contexts)?
OP9: Normative specification and governance.
Who chooses the axiom mix, loss weights, and robustness threat model when the rule is learned, and how can these choices be made transparent and contestable?
OP10: Robustness to strategic coalitions with realistic constraints.
Can adversarial training be grounded in social-choice-native threat models (coalition size, limited information, bounded rationality) rather than generic perturbation norms?
OP11: Interpretable structure discovery.
Can we reliably extract human-legible descriptions (e.g., scoring vectors, pairwise tournaments, committee scoring rules) from trained networks so that the learned rule can be debated and audited like a classical rule?
OP12: Bridging single-winner, multi-winner, and continuous outcomes.
Is there a unified differentiable framework that scales from single-winner elections to committees, rankings, and continuous allocations while preserving the relevant axioms in each regime?
4 Participatory Budgeting and Complex Allocation
Table 3: Comparison of classical participatory budgeting rules and differentiable or learned PB pipelines. The table highlights trade-offs between representation guarantees, objective flexibility, feasibility enforcement, computational scalability, and auditability.
Rule / Pipeline Ballot Proportionality Objective Feasible Exhaustive Scalable Auditability
Greedy (approval) Peters et al. (2021) Approval None Utilitarian High High
Greedy (score-to-cost) Brandt et al. (2016) Approval / Utility None Utilitarian High High
Phragmén-style methods Aziz et al. (2018); Brandt et al. (2016) Approval PJR (approval) Load balancing Medium High
Method of Equal Shares (MES) Peters et al. (2021) Approval / Utility EJR-style Proportionality-first Medium High
Equal-shares completion heuristics Kraiczy et al. (2024; 2025) Approval / Utility EJR-style Proportionality-first Low–Medium Medium
Adaptive MES (AMES) Kraiczy et al. (2024) Approval / Utility EJR-style Proportionality-first High High
Learned PB (rule mimicry) Fairstein et al. (2024) Approval / Utility Inherited Target rule High Medium
Learned PB (decision-focused) Fairstein et al. (2024); Amos & Kolter (2017); Wilder et al. (2019) Approval / Utility Soft constraints Multi-objective High Medium
LLM-based PB (text $\rightarrow$ allocation) Damle & Faltings (2025) Natural language Implicit / heuristic Prompt-dependent High Low
LLM + certified solver Damle & Faltings (2025); Amos & Kolter (2017) Text $\rightarrow$ structured Solver-dependent Hybrid Medium Medium–High
Notes. “Proportionality” refers to representation guarantees such as JR/EJR for approval PB and equal-shares-style guarantees for cost-aware settings, see Sec 4.1. “Feasible” indicates satisfaction of the budget constraint by construction. Entries marked require either solver-layer enforcement or post-processing. Scalability refers to practical deployment at municipal scale; learned methods typically incur higher training cost but low inference cost.
Participatory budgeting (PB) extends social choice from selecting a single winner to selecting a feasible portfolio of projects subject to a hard budget constraint. This seemingly modest change transforms aggregation into a knapsack-style combinatorial optimisation problem coupled with normative requirements such as proportionality, minority representation, and interpretability Peters et al. (2021); Brandt et al. (2016). In practice, PB deployments are now routine across municipalities and institutions, yet organisers face an enduring design dilemma: different aggregation rules induce materially different distributions of benefits, and there is rarely consensus on which normative trade-off is appropriate Peters et al. (2021); Kraiczy et al. (2024).
Differentiable approaches reposition this dilemma as a learning problem. Rather than committing ex ante to a named rule (e.g., greedy approval knapsack, Phragmén-style methods, or the Method of Equal Shares), a designer specifies target metrics and trains a model to map ballots (and increasingly, deliberative text) to feasible portfolios Fairstein et al. (2024); Damle & Faltings (2025). This chapter reviews (i) classical PB axioms and rules that serve as the normative and algorithmic baseline, (ii) supervised and differentiable PB as “learned aggregation over portfolios”, (iii) strategic and multi-round PB through reinforcement learning, (iv) the emergence of LLMs as preference-extraction and allocation agents, and (v) how PB connects to a broader landscape of complex allocation problems—committee selection, constrained public decision-making, and fair division—where differentiability offers new tools but also new failure modes.
Table 3 summarizes classical and learned participatory budgeting pipelines, highlighting which properties (representation, feasibility, and auditability) are guaranteed by design versus only evaluated empirically.
4.1 The PB model: ballots, feasibility, and normative desiderata
From preference aggregation to constrained optimisation.
A standard PB instance consists of a set of projects $P=\{1,\dots,m\}$ with costs $c_{j}\geq 0$, a total budget $B$, and voters $i\in\{1,\dots,n\}$ who express preferences over projects via approval, rankings, or cardinal utilities. The outcome is a funded set $W\subseteq P$ such that $\sum_{j\in W}c_{j}\leq B$. With approval ballots, the simplest objective is to maximise the number of satisfied approvals subject to feasibility (a maximum-coverage variant of knapsack), which is NP-hard in general and motivates heuristics Peters et al. (2021); Brandt et al. (2016). With cardinal utilities, one might maximise utilitarian welfare $\sum_{i}u_{i}(W)$ or Nash social welfare $\prod_{i}(u_{i}(W)+\epsilon)$, each encoding a different normative stance on inequality and trade-offs Caragiannis et al. (2016); Cole & Gkatzelis (2015).
Axioms: proportionality and representation in PB.
PB is often motivated as a proportionality problem: minority groups should not be systematically excluded from funding. Axioms adapted from multi-winner voting—such as justified representation (JR) and extended justified representation (EJR)—formalise these guarantees when voters submit approvals Peters et al. (2021); Aziz et al. (2018); Brandt et al. (2016). The central difficulty is that proportionality interacts with feasibility: high-cost projects can crowd out representation even when they have broad support, and conversely a proportional outcome may be inefficient under utilitarian metrics Peters et al. (2021). This tension is precisely where modern PB theory has focused: identifying rules that are polynomial-time computable, satisfy strong proportionality axioms, and still provide efficiency guarantees.
Proportionality notions in approval-based PB.
In approval-based participatory budgeting, proportional representation is commonly formalized using a hierarchy of axioms adapted from multi-winner voting. Justified Representation (JR) requires that any sufficiently large and cohesive group of voters is represented by at least one approved project. Proportional Justified Representation (PJR) strengthens this requirement by demanding representation proportional to group size. Extended Justified Representation (EJR) is stronger still, requiring that some voter in each such group individually receives their proportional share of funded projects. These axioms satisfy a strict implication hierarchy, $\text{EJR}\Rightarrow\text{PJR}\Rightarrow\text{JR}$, and serve as the main normative benchmarks for approval-based PB rules Peters et al. (2021); Brandt et al. (2016).
4.2 Classical PB rules: from greedy knapsack to equal-shares proportionality
Greedy and knapsack-style heuristics.
Many real-world PB deployments historically default to greedy variants: repeatedly fund the most popular affordable project (or the project with best score-to-cost ratio) until the budget is exhausted. These rules are appealing for transparency and simplicity but can be majoritarian, systematically under-serving cohesive minorities, especially when project costs vary widely Peters et al. (2021). Their behaviour is tightly linked to the geometry of costs: when many small projects exist, greedy can approximate broad satisfaction; when costs are skewed, greedy can concentrate spending.
Proportional PB with additive utilities.
A foundational modern treatment is Proportional Participatory Budgeting with Additive Utilities Peters et al. (2021), which formalises PB as a generalisation of approval-based multi-winner voting and studies rules that satisfy proportionality while remaining computationally feasible. This work crystallises a design pattern: interpret the budget as distributed purchasing power across voters, and fund projects by charging supporters in a way that respects equal treatment and proportional representation. It also clarifies when and how proportionality axioms extend beyond approvals to additive utilities.
The Method of Equal Shares (MES): fairness via cost-sharing.
Among proportional PB rules, the Method of Equal Shares (MES) has emerged as a widely studied and practically deployed standard. Conceptually, MES assigns each voter an equal virtual budget share ($B/n$) and funds projects by charging their supporters, selecting affordable projects according to a rule that balances marginal cost against support Peters et al. (2021). MES satisfies demanding proportionality guarantees (in particular, EJR in approval settings under appropriate formulations) and embodies a concrete ethical principle: spending should be “paid for” by those who benefit, but everyone starts with equal purchasing power Peters et al. (2021).
Exhaustiveness and practical completion heuristics.
A known drawback is that MES can be non-exhaustive: it may leave budget unused even though additional projects could be funded without violating the constraint Kraiczy et al. (2024). This has led to completion heuristics in practice (e.g., “add-one” variants that inflate a virtual budget until the funded set becomes maximal), which can be computationally inefficient and conceptually unclean. Recent theoretical work proposes algorithmic refinements that make equal-shares methods more practical at scale. For instance, Adaptive MES (AMES) exploits similarities across nearby budget levels and provides a certifiable proportionality check, turning proportionality from an abstract guarantee into an auditable artefact Kraiczy et al. (2024). Related work studies efficient completion procedures (e.g., add-opt) for equal-shares variants to avoid repeated re-computation Kraiczy et al. (2025). These developments are crucial baselines for differentiable PB: learned mechanisms must be compared not only to classical MES, but to its algorithmically improved, deployable descendants.
Efficiency guarantees for proportional rules.
A recurring concern is whether proportionality sacrifices too much welfare. “Utilitarian Guarantees for the Method of Equal Shares” (and related analyses) study bounds on the utilitarian performance of MES, showing that proportional methods can enjoy strong welfare guarantees that are asymptotically tight among proportional rules Baychkov et al. (2025). Such results matter for learning-based PB in two ways: (i) they set a principled target for “efficiency without abandoning proportionality”, and (ii) they provide benchmark curves that learned rules should be expected to match or exceed only by relaxing (or reinterpreting) proportionality constraints.
4.3 Learning aggregation rules in PB: supervised and differentiable pipelines
PB as a learnable mapping from approval profiles to feasible portfolios.
Learning Aggregation Rules in Participatory Budgeting: A Data-Driven Approach introduces a general framework for learning PB rules from data Fairstein et al. (2024). The central object is a model $f_{\theta}$ that maps an approval profile (or richer features) to a feasible funded set, trained to optimise a designer-specified objective that can blend efficiency and fairness. In contrast to “mimicry” of a known rule, the contribution is the ability to synthesise compromise rules: the organiser writes down a loss (e.g., total approvals satisfied, representation proxies, inequality penalties such as a Gini coefficient over voter satisfaction), and the learned rule implements that compromise Fairstein et al. (2024).
Architectural and algorithmic choices: feasibility as a first-class constraint.
Learning PB differs from learning single-winner voting rules because feasibility is global and combinatorial. Common strategies include: (i) continuous relaxations that output fractional selections followed by rounding, (ii) differentiable projections onto knapsack-like polytopes, and (iii) optimisation layers that embed a constrained solver inside the network and backpropagate through it via implicit differentiation Amos & Kolter (2017). The broader differentiable-optimisation literature provides tools for these pipelines: OptNet for differentiable quadratic programs Amos & Kolter (2017), Sinkhorn-based relaxations for permutations and matchings Mena et al. (2018), and more generally, a family of “decision-focused learning” methods that train predictors end-to-end against downstream combinatorial objectives rather than surrogate prediction losses Wilder et al. (2019).
In PB, the design problem is not merely selecting an approximation method, but aligning it with normative requirements. A smooth relaxation may unintentionally bias toward “spread-out” fractional solutions that, after rounding, favour low-cost projects; a hard projection may preserve feasibility but introduce non-smoothness that destabilises training. Consequently, PB has become a natural testbed for a more general question: when should we enforce constraints by architecture (solver layers), by penalties (Lagrangians), or by post-processing (rounding), and what is the normative cost of each choice?
Generalisation across instance size and electorate shift.
A salient empirical claim in learned PB is cross-scale generalisation: models trained on small synthetic PB instances can generalise to larger, real PB elections Fairstein et al. (2024). This is non-trivial because the output space grows exponentially with the number of projects, and because real PB data often exhibits structured correlations (neighbourhoods, demographics, and thematic clusters). Set- and graph-based architectures (e.g., DeepSets and attention mechanisms) are natural candidates for such generalisation because they encode permutation invariances and can operate on variable-sized inputs Zaheer et al. (2017); Lee et al. (2019). However, generalisation in PB is also fundamentally normative: a rule that generalises in terms of objective value may not generalise in terms of proportionality or minority protection under distribution shift, motivating audit protocols analogous to those developed for neural voting rules.
4.4 Strategic and multi-round PB: reinforcement learning and decision support
From static aggregation to interactive compromise.
Classical PB models treat ballots as fixed. In many deployments, however, PB is deliberative: voters discuss, revise preferences, and coordinate. Fair Compromises in Participatory Budgeting: a Multi-Agent Deep Reinforcement Learning Approach models PB as an interactive process in which agents adapt their support across rounds, learning strategies that lead to collectively fairer outcomes Adams et al. (2025). A key empirical finding is that compromise may emerge through coordination on lower-cost projects that deliver broad satisfaction and representation, rather than conflict over expensive flagship proposals Adams et al. (2025).
Interpretation: PB as mechanism-mediated coordination.
This line reframes PB from a single-shot social choice rule into a coordination mechanism that can steer the system toward Pareto-improving equilibria. It also surfaces a conceptual split:
  • Aggregation-as-decision: compute the funded set from reported preferences.
  • Aggregation-as-process: design interaction protocols (recommendations, negotiation aids, explanations) that help the electorate reach better compromises.
Reinforcement learning is particularly well suited to the second view because it can optimise policies over trajectories (discussion rounds, proposal revisions) and evaluate long-horizon fairness criteria. Yet it also inherits the governance challenges of learned mechanisms: policy behaviour can be sensitive to reward shaping, may exploit proxy metrics, and is difficult to certify.
4.5 LLMs for PB and resource allocation: from unstructured preferences to feasible budgets
PB as a benchmark for language-to-allocation reasoning.
PB increasingly occurs in online environments where preferences are expressed in natural language—comments, discussions, and proposals—rather than clean approval vectors. LLMs for Resource Allocation: A Participatory Budgeting Approach to Inferring Preferences treats PB as both a practical application and an evaluative benchmark for LLM reasoning under feasibility constraints Damle & Faltings (2025). The work separates two capabilities: extracting structured preferences from text (a strength of modern LLMs) and solving the combinatorial allocation under budget constraints (more fragile and prompt-dependent) Damle & Faltings (2025).
Prompting as mechanism design.
The study compares prompting strategies that mirror classical algorithms: “greedy” selection, explicit knapsack optimisation framing, and iterative hill-climbing refinement Damle & Faltings (2025). This illustrates a broader theme in differentiable social choice: when the “aggregator” is a foundation model, the mechanism is partly specified by prompts, scaffolds, and tool use. The normative and technical questions then shift:
  • Does the LLM respect proportionality, or does it default to majority heuristics?
  • Can we make feasibility guaranteed (e.g., via tool-augmented solvers) rather than hoped for?
  • How do we audit and reproduce outcomes when the mechanism includes stochastic generation?
These issues motivate hybrid designs: LLMs for preference inference and explanation, coupled with certified optimisation layers (or classical PB rules) for the final allocation.
4.6 Beyond PB: complex allocation as differentiable social choice
Participatory budgeting is one point in a larger design space of complex collective allocation problems. The key commonality is that the social choice output is structured (a set, ranking, matching, schedule, or flow) and constrained (budgets, capacities, diversity requirements, fairness constraints). Differentiability provides a unifying engineering language—optimise an objective subject to constraints—but the normative content depends on what objective and constraints are chosen.
Committee selection and multi-winner voting with costs.
PB can be viewed as multi-winner voting with heterogeneous costs rather than unit-cost candidates. Many representation axioms (JR/EJR) and algorithmic ideas (Phragmén methods, priceability) originate in approval-based committee selection and carry over to PB Aziz et al. (2018); brandl2016handbook. Equal-shares methods can be interpreted as cost-aware generalisations of proportional committee rules, reinforcing the conceptual continuity between PB and multi-winner elections Peters et al. (2021).
Fair division and welfare objectives under constraints.
At a deeper level, PB is also an instance of fair division of public goods: allocate a shared budget across projects to maximise welfare while respecting equity. This connects to the literature on Nash social welfare (NSW), which provides a principled compromise between efficiency and fairness and has become influential in allocating indivisible goods Caragiannis et al. (2016); Cole & Gkatzelis (2015). Differentiable approaches can operationalise NSW-style objectives in PB—potentially via smooth approximations or log-sum formulations—while embedding feasibility through solver layers Amos & Kolter (2017); Wilder et al. (2019). However, the fairness notions differ: PB representation axioms are group-based and coalition-sensitive, while NSW fairness is typically individual-centric. Designing objectives that respect both perspectives remains non-trivial.
Differentiable combinatorial optimisation: relaxations, solvers, and decision-focused learning.
Complex allocation problems have motivated a broad ML toolkit: (i) relax-and-round methods (continuous surrogates then discrete rounding), (ii) implicit differentiation through optimisation problems (e.g., OptNet) Amos & Kolter (2017), (iii) continuous approximations to discrete structures such as matchings and permutations via Sinkhorn operators and Gumbel perturbations Mena et al. (2018), and (iv) decision-focused learning, which trains predictors directly for downstream decision quality Wilder et al. (2019). PB sits at an attractive intersection where these tools can be evaluated against socially meaningful metrics (representation, inequality, and transparency), rather than purely operational cost.
Auditing and legitimacy in learned allocation.
Compared to auctions (where IC is central) and voting (where anonymity/neutrality dominate), PB places proportionality and budget feasibility at the forefront. Learned PB therefore inherits two distinct audit burdens:
  1. 1.
    Constraint correctness: Is the funded set feasible, and does it satisfy promised proportionality bounds (or their proxies)?
  2. 2.
    Normative faithfulness: Does the learned compromise reflect the organiser’s stated trade-offs, or has the model exploited proxy objectives and distributional quirks?
The second is particularly subtle: a model can optimise a fairness proxy (e.g., satisfaction Gini) while systematically failing coalition-based representation axioms that matter for democratic legitimacy Peters et al. (2021). This motivates “axiomatic evaluation suites” for PB analogous to those developed for neural voting rules, but adapted to cost constraints and portfolio outputs.
4.7 Open Problems
OP13: Hard proportionality constraints in learned PB.
Can we design architectures or solver-in-the-loop training procedures that enforce PB proportionality guarantees (e.g., EJR-style properties for approvals) as hard constraints rather than soft penalties?
OP14: Generalisation of fairness under distribution shift.
How can learned PB rules preserve representation guarantees when the electorate changes, project costs shift, or new categories of projects appear (i.e., under real municipal non-stationarity)?
OP15: Transparent compromise specification.
What is the right interface for organisers to specify a normative objective (fairness–efficiency trade-offs, minority protections, geographic constraints) such that the resulting learned rule is explainable, contestable, and reproducible?
OP16: LLM-based PB with certified feasibility.
Can we combine LLM preference inference and deliberation support with optimisation layers or classical PB rules to guarantee feasibility and provide verifiable proportionality while still benefiting from language understanding?
OP17: Strategic behaviour and long-horizon dynamics.
When PB is deliberative and repeated, how should we model strategic adaptation, coalition formation, and agenda-setting, and can MARL-based decision support avoid steering the process toward manipulative equilibria?
OP18: Unified benchmarks for complex allocation mechanisms.
Can the community develop standardised benchmarks (datasets, axiomatic test suites, and threat models) that allow meaningful comparison across classical PB rules, learned aggregators, MARL processes, and LLM-based allocators?
5 Liquid Democracy and Decentralized Aggregation
Liquid democracy (LD), also known as delegable proxy voting, sits between direct and representative democracy: each voter may either vote directly or delegate their vote to another voter, who may further re-delegate, so that voting power flows along directed edges in a delegation graph. The system is conceptually simple yet structurally rich: transitivity induces power concentration, delegation cycles, and path-dependent notions of representation. This makes LD an unusually natural bridge between social choice and distributed machine learning. In federated learning (FL), clients compute local updates and an aggregator combines them into a global model; the choice of who communicates, whose update is trusted, and how influence propagates is—mathematically and normatively—a problem of aggregation on graphs McMahan et al. (2017); Kesari & Kesari (2025).
This chapter reviews LD through two intertwined lenses. The first is the classical computational social choice perspective: models of delegation, cycles, strategic incentives, power indices, and epistemic performance (truth-tracking). The second is the “differentiable social choice” perspective, where delegation architectures are repurposed to design communication-efficient, robust, and decentralized aggregation protocols for FL and edge networks. Across both, a recurring theme is that the delegation graph is not merely an input—it is a design variable. Learning, optimizing, or adapting delegation structures raises new questions about governance, robustness, and verifiability.
5.1 Delegable proxy voting as a social choice primitive
Basic models and resolution of delegation chains.
The canonical LD setting is binary voting (two alternatives), where each voter either casts a direct vote or delegates to a proxy, yielding a directed graph in which weights accumulate along paths ending at casting voters Christoff & Grossi (2017); Kahng et al. (2021). Practical systems must resolve three technical choices that strongly shape outcomes: (i) what happens on delegation cycles (e.g., $i\!\to\!j\!\to\!i$), (ii) whether delegation is issue-by-issue or persistent across issues, and (iii) what information voters use when selecting delegates (local neighbor information versus non-local global properties) Christoff & Grossi (2017); Kahng et al. (2021).
Cycles are not an edge case.
Cycles are often treated as a pathology, but they are structurally unavoidable in large graphs unless the design forbids them by construction. Christoff and Grossi highlight cycles as a central conceptual obstacle and study how cycles interact with rationality when issues are logically related (binary aggregation with integrity constraints), where delegation can propagate inconsistencies Christoff & Grossi (2017). More recent work reframes cycles as an equilibrium phenomenon: in default delegation models (long-run delegation decisions under uncertainty), cycles can emerge among rational agents, yielding a game-theoretic justification rather than a mere implementation bug Grossi & Nitsche (2025).
Expressivity of delegations.
Realistic delegation preferences often exceed “delegate to exactly one person.” Richer models allow ranked delegates, partial delegations, or topic-dependent delegation, which increases expressivity but raises computational questions. Recent complexity analyses study how expressive delegation languages change the tractability of computing outcomes and verifying desirable properties Tyrovolas et al. (2024). These results matter for differentiable approaches: a learned aggregator that implicitly implements fractional or multi-proxy delegation may be powerful, but also harder to audit.
5.2 Strategic incentives and endogenous delegation
Delegation as a game.
LD is often motivated as “vote with your trust,” but in strategic settings trust is endogenous. Game-theoretic models study equilibria where agents choose whether to vote or delegate based on expected competence, effort, or influence concerns. “On Rational Delegations in Liquid Democracy” formalizes such a delegation game, analyzing equilibrium structure and the interplay between competence assumptions and delegation outcomes Bloembergen et al. (2019). A complementary line asks how the desire for power (rather than correctness) affects delegation choices. Recent work develops power indices for delegable proxy systems and studies equilibria under power-seeking behavior, illustrating how LD can drift toward oligarchic concentration even absent explicit manipulation Zhang & Grossi (2021).
Influence and concentration as design targets.
The tension between epistemic benefits (delegating to experts) and democratic concerns (power concentration) has motivated formal influence measures and constraints. “Rethinking Influence in Liquid Democracy” argues that static delegation analysis misses the dynamic nature of influence accumulation and studies how influence should be measured and controlled over time Grossi & Nitsche (2025). This perspective aligns closely with FL, where influence concentration can correspond to over-weighting a small set of clients, risking bias and poisoning vulnerability.
5.3 Epistemic performance: when does liquid delegation track truth?
The epistemic promise and its limits.
A standard normative justification for LD is epistemic: if some voters are more competent, delegation can concentrate weight on experts and improve the probability of choosing the correct outcome Berinsky et al. (2023). However, algorithmic analysis shows that this is not automatic. Kahng et al. model an electorate with heterogeneous competence and ask whether local delegation mechanisms can be guaranteed to outperform direct majority voting; they provide negative results for local-only rules and propose non-local mechanisms that can provably improve truth-tracking under mild assumptions Kahng et al. (2021).
Armstrong and Larson sharpen the critique in “On the Limited Applicability of Liquid Democracy,” arguing that the epistemic case depends on restrictive conditions (e.g., a single objectively correct outcome and competence signals aligned with delegation) and that LD can fail to deliver benefits in broader settings Armstrong & Larson (2021). This cautions against treating LD as a universal improvement over direct voting; instead, its performance is distribution- and context-dependent, echoing the broader shift from worst-case axioms to data-aware evaluation in differentiable social choice.
New evidence and a “high-probability” view.
“Tracking Truth with Liquid Democracy” develops a theoretical model and extensive experiments (including human-subject evidence) showing that liquid delegation can typically satisfy probabilistic variants of “positive gain” and “do no harm” under minimal assumptions, pushing back against overly pessimistic interpretations of earlier worst-case analyses Berinsky et al. (2023). A complementary mechanistic refinement is to allow weighted delegations: Zhang and Grossi show that weighting proxies can improve group accuracy in certain regimes, and that equilibria with weighted delegations are never worse (and sometimes better) than standard delegations Zhang & Grossi (2021).
Together, these works suggest a more nuanced message suitable for “differentiable social choice”: LD is best viewed as a family of aggregation architectures whose epistemic performance can be optimized for a population and a context, rather than as a single rule with uniform guarantees.
5.4 Costs, feasibility, and control: the resource realism of delegation
Delegation under budget constraints.
In practice, voting and delegating are costly: cognitive effort, time, attention, or (in digital systems) communication bandwidth. “The Cost Perspective of Liquid Democracy: Feasibility and Control” introduces an approval-based LD model with explicit budget constraints on voting and delegating costs, and studies the problem of selecting a set of casting voters such that every voter is represented by a path to a caster while minimizing total cost Alouf-Heffetz et al. (2025). This formulation reveals LD as a constrained graph-coverage problem with a distinctive normative requirement: universal representation under resource limits.
Complexity and vulnerability to manipulation via system parameters.
Alouf-Heffetz et al. further analyze computational complexity variants (e.g., controlling delegation path lengths and limiting power concentration) and study strategic control: whether an external actor can manipulate budget parameters or components of the system to force centralization or alter power distributions Alouf-Heffetz et al. (2025). This is an important bridge to distributed ML: if delegation-like architectures are used to aggregate gradients, then “controlling costs” corresponds to communication scheduling and topology constraints, and “strategic control” corresponds to adversarial topology or incentive manipulation.
5.5 From liquid voting to federated learning: fluid delegation as decentralized aggregation
Federated aggregation as a voting rule.
The dominant FL baseline, FedAvg, aggregates client updates via a weighted mean, typically proportional to local data size McMahan et al. (2017). In the social choice analogy, this is “one-person-one-vote” (or one-datum-one-vote): every participating client transmits to a central server each round, and influence is a linear function of dataset size. This design is bandwidth-heavy and vulnerable to poisoning and non-i.i.d. heterogeneity, motivating alternatives that select or reweight clients, use robust aggregation (median/trimmed mean, Krum), or move toward decentralized (gossip-based) protocols Blanchard et al. (2017); Bonawitz et al. (2017).
Viscous-retained democracy (FedVRD): delegation for bandwidth and robustness.
“Fluid Democracy in Federated Data Aggregation” explicitly maps LD topology onto FL: clients evaluate neighbor competence (e.g., via local validation loss) and delegate their model updates to more competent peers in a peer-to-peer graph, reducing direct communication with the server Kesari & Kesari (2025). The proposed protocol, viscous-retained democracy (FedVRD), introduces a “viscosity” parameter that dampens influence flow, aiming to preserve epistemic gains (routing updates through competent nodes) without allowing unbounded influence accumulation Kesari & Kesari (2025). The paper emphasizes two practical payoffs: (i) communication savings by reducing the number of server-bound uploads, and (ii) improved robustness to poisoning, as viscous delegation can dilute adversarial influence and limit amplification through delegation chains Kesari & Kesari (2025).
Relation to robust and secure aggregation.
FedVRD is complementary to robust aggregation rules developed for Byzantine-resilient FL. Classical defenses include Krum and its variants, which select updates close to the majority in Euclidean space Blanchard et al. (2017), trimmed mean/coordinate-wise median aggregators, and cryptographic secure aggregation protocols that hide individual updates from the server Bonawitz et al. (2017). The LD-inspired shift is architectural: rather than relying purely on server-side filtering, delegation places part of the aggregation burden on the network, enabling hierarchical or multi-hop aggregation. This resembles communication-efficient FL designs (client selection, hierarchical FL) but adds a trust graph as a principled organizing primitive Kesari & Kesari (2025).
5.6 Learning and optimizing delegation graphs: toward differentiable decentralization
Delegation as a learnable routing policy.
Once delegation is viewed as a routing policy for votes or gradients, it becomes amenable to learning. The objective can be epistemic accuracy (truth tracking), robustness (poisoning resistance), fairness (limiting influence concentration), or cost (communication/effort) Kahng et al. (2021); Alouf-Heffetz et al. (2025); Kesari & Kesari (2025). In FL, one can imagine optimizing delegation edges, viscosity parameters, or delegate selection thresholds end-to-end against validation accuracy and communication budgets. In civic LD, one can imagine optimizing delegation recommendation systems under constraints intended to preserve autonomy and avoid manipulative concentration.
Topology dependence and the governance of graphs.
Both LD and FL variants are highly topology-dependent: who can delegate to whom (or which clients are neighbors) determines both performance and vulnerability. This raises governance questions that are easy to miss in purely algorithmic treatments: who defines the graph, how dynamic rewiring is controlled, and how to prevent graph-level attacks (Sybil insertions, bridge capture, or collusion among central nodes). The cost-and-control perspective formalizes parts of this via strategic control models Alouf-Heffetz et al. (2025); FL security literature offers threat models and defenses, but integrating them with delegation semantics remains an open area Blanchard et al. (2017); Bonawitz et al. (2017).
5.7 Open Problems
OP19: Certified limits on influence concentration.
Can we design liquid (or fluid) aggregation rules with hard, verifiable bounds on influence accumulation that remain effective under realistic, dynamic delegation graphs?
OP20: Delegation under adversarial topology.
How should delegation protocols be redesigned when the underlying trust/communication graph is partially controlled by an adversary (e.g., Sybil nodes, bridge capture, collusion), and what threat models are normatively appropriate?
OP21: End-to-end learning of delegation with guarantees.
Can we learn delegation policies (edges, weights, viscosity) end-to-end for accuracy and bandwidth while providing post-training certificates for robustness, representation, and bounded regret?
OP22: Cycles as a feature, not a bug.
When delegation cycles emerge as rational equilibria, how should systems resolve them in a way that preserves legitimacy and avoids pathological deadlocks or power anomalies?
OP23: Interfacing liquid delegation with secure aggregation.
Can LD-inspired multi-hop aggregation be combined with cryptographic secure aggregation so that privacy is preserved while competence-based delegation still yields epistemic and robustness gains?
OP24: Legitimacy of learned or recommended delegations.
If platforms recommend delegates (or FL clients auto-select delegates based on “competence”), how do we ensure transparency and prevent subtle steering toward centralized control or biased outcomes?
6 AI Alignment as a Social Choice Problem
6.1 Section overview: the hidden ballot in the loss function
Modern alignment pipelines for large language models (LLMs) are, at their core, preference aggregation systems. Reinforcement Learning from Human Feedback (RLHF) learns from many pairwise comparisons—“output $A$ is better than output BB”—collected from annotators who may disagree for principled, contextual, or idiosyncratic reasons Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022). Direct Preference Optimization (DPO) and related “direct alignment” objectives bypass explicit reward modeling but still optimize a loss induced by preference data Rafailov et al. (2023). In both cases, the training procedure compresses heterogeneous human judgments into a single reward model or policy, thereby performing an implicit social choice operation.
Recent theoretical work makes this connection explicit and consequential. Distributional Preference Learning (DPL) proves that standard Bradley–Terry–Luce (BTL) preference modeling—a workhorse in RLHF—implicitly aggregates across heterogeneous annotator contexts via the Borda count Siththaranjan et al. (2024). This equivalence is not merely rhetorical: it imports classical social choice pathologies (e.g., indifference-to-majority phenomena and clone vulnerabilities) into alignment, and it reframes “reward hacking” and “jailbreaks” as failures of aggregation under hidden context Siththaranjan et al. (2024). In parallel, Conitzer et al. argue that social choice theory should guide alignment precisely because disagreement is irreducible: the problem is not to find a single “true” preference, but to choose a principled way to aggregate diverse feedback into system behavior Conitzer et al. (2024). Dai and Fleisig sharpen the mapping by emphasizing where RLHF differs from canonical voting settings (e.g., strategic feedback collection, non-stationary policies, and distribution shift), cautioning against naive transfer of theorems while endorsing social-choice-native desiderata as analytical tools Dai & Fleisig (2024).
This chapter reviews alignment as a social choice problem across five themes: (i) the social choice structure of RLHF and direct alignment objectives, (ii) heterogeneity-aware learning (distributional and pluralistic alignment), (iii) axiomatic and welfare-theoretic perspectives on what “collective preference” should mean, (iv) strategic behavior, manipulation, and mechanism design for feedback pipelines, and (v) governance and decentralization: from “one reward model” to constitutional and participatory alignment.
Method Loss / Objective Setting Corresp. Voting Rule Axioms / Guarantees Diff.
Majority
resp.
Pareto
IIA/LIIA
Condorcet
(A) Standard alignment objectives as implicit voting rules
RLHF reward model (BTL / logistic) $\sum_{(x,y^{+},y^{-})}\log\!\Big(1+\exp\big(r_{\theta}(x,y^{-})-r_{\theta}(x,y^{+})\big)\Big)$ Pairwise preferences (RLHF) Borda-like scoring
Exponential pairwise loss $\sum_{(x,y^{+},y^{-})}\exp\big(r_{\theta}(x,y^{-})-r_{\theta}(x,y^{+})\big)$ Pairwise preferences (RLHF-style) Borda-like scoring
Hinge / margin ranking loss $\sum_{(x,y^{+},y^{-})}\max\!\Big(0,\,1+r_{\theta}(x,y^{-})-r_{\theta}(x,y^{+})\Big)$ Pairwise preferences (ranking / RLHF-style) Borda-like scoring
Direct Preference Optimization (DPO) $\mathbb{E}_{(x,y^{+},y^{-})}\!\Big[-\log\sigma\!\big(\beta(\log\pi_{\theta}(y^{+}|x)-\log\pi_{\theta}(y^{-}|x))\big)\Big]$ Direct policy optimization Tournament solution (loss-induced) var. var. var.
(B) Heterogeneity-aware / pluralistic alignment (changes the “electorate”)
Distributional Preference Learning (DPL) Latent-context model; preserves heterogeneity in preference signal (beyond a single scalar reward) Hidden-context preferences Borda-like aggregation in the BTL baseline
Pluralistic Alignment Framework (PAL) Latent-variable / ideal-point style preference model; user- or subgroup-conditioned reward/policy Mixture of preference types Mixture/committee-style aggregation type-dep. type-dep. type-dep.
(C) Axiomatic loss design (Differential Voting)
Soft Copeland $y\,s_{\tau,\beta}(\Delta)+\tfrac{\lambda}{2}\Delta^{2}$ Axiomatic loss for pairwise aggregation Copeland (smooth surrogate)
Soft Kemeny $\sigma(-y\,\Delta/\tau)$ Axiomatic loss for pairwise aggregation Kemeny (smooth Kendall surrogate)
Majority-based objectives (family) Any objective depending only on pairwise majority directions Pairwise-majority tournament (Pairwise) Majority aggregation
LCPO (leximax Copeland subj. to PO) Leximax Copeland with Pareto constraint (implemented via LP feasibility checks) Hard-constraint rule layer Copeland (PO-constrained leximax)
Table 4: Alignment objectives as implicit social choice rules. The choice of loss in RLHF-style preference learning (and related direct alignment objectives) instantiates a concrete aggregation principle with specific axiomatic trade-offs. BTL-style objectives correspond to Borda-like scoring under heterogeneous/hidden contexts Siththaranjan et al. (2024), inheriting known violations of majority responsiveness and Condorcet consistency Brandt et al. (2016); Ge et al. (2024). DPO can be interpreted as optimizing a loss-induced tournament solution over sampled candidates, so axiom satisfaction is generally sampling- and protocol-dependent Rafailov et al. (2023); Dai & Fleisig (2024); Conitzer et al. (2024). Differential Voting constructs differentiable surrogates for Condorcet-style rules (e.g., Copeland, Kemeny), making the intended aggregation principle explicit at the loss level An et al. (2026).
6.2 RLHF as implicit voting: BTL, Borda, and the geometry of preference learning
From pairwise labels to a social ranking.
RLHF typically fits a reward model $r_{\theta}(x,y)$ so that, for a prompt $x$ and two candidate outputs $y^{+}$ and $y^{-}$,
$\Pr(y^{+}\succ y^{-}\mid x)\approx\sigma\!\big(r_{\theta}(x,y^{+})-r_{\theta}(x,y^{-})\big),$ (2)
which is the BTL likelihood used widely in preference learning Christiano et al. (2017); Stiennon et al. (2020). When labels come from many annotators and contexts, the learning objective averages log-likelihood contributions across the dataset; the critical question is: what aggregation principle does this averaging implement?
DPL answers this precisely. Modeling “hidden context” as a latent variable indexing annotator type, rubric, or situational factors, Siththaranjan et al. show that minimizing the standard BTL negative log-likelihood implicitly aggregates these latent contexts according to Borda count Siththaranjan et al. (2024). In classical social choice terms, Borda assigns each alternative a score equal to its average rank, favoring consensus candidates and smoothing over sharp disagreement Young (1974); Saari (1994). In alignment terms, this predicts a familiar empirical outcome: models that feel like “the safe average” can arise not (only) from explicit safety tuning, but from the statistical structure of the loss.
Alignment objectives as social choice rules.
A central message of this section is that modern alignment pipelines are not value-neutral optimization procedures: every preference-learning or direct-alignment objective implicitly defines a collective decision rule over candidate model behaviors. Table 4 makes this connection explicit by mapping commonly used alignment losses to their corresponding social choice interpretations and axiomatic properties. In particular, widely used RLHF-style objectives based on Bradley–Terry–Luce likelihoods implement Borda-like aggregation under heterogeneous or hidden contexts, inheriting known violations of majority responsiveness and Condorcet consistency. Pluralistic approaches modify the structure of the electorate but do not, by themselves, resolve aggregation pathologies. By contrast, Differential Voting specifies the aggregation principle directly at the level of the loss, enabling differentiable surrogates for Condorcet-consistent rules and making normative trade-offs explicit and auditable.
Social choice pathologies become alignment failure modes.
The Borda lens imports a mature theory of failure. Borda can select an outcome that is not a Condorcet winner (an option that beats every other option head-to-head) de Condorcet (1785). It is susceptible to clone effects, where adding similar options can change the winner, and it can underweight intense majority preference in favor of broad second-choice acceptability Tideman (2006). Translating to RLHF, one expects: (i) majority dissatisfaction when a large subgroup strongly prefers a style or value that is diluted by averaging, (ii) minority erasure when rare but important preferences are smoothed away, and (iii) strategic vulnerabilities because if contributors understand the aggregation rule, they can shape outcomes by how they label and which comparisons they provide Siththaranjan et al. (2024).
Dai and Fleisig emphasize an additional geometric mismatch: in voting, candidates are discrete; in RLHF, outputs lie on a vast, structured manifold and the model changes as data is collected Dai & Fleisig (2024). This makes classical impossibilities (e.g., Arrow-style) suggestive but not directly dispositive; nevertheless, they remain valuable as warnings that no aggregation rule will satisfy all desiderata simultaneously Arrow (2020); Sen (1986).
6.3 From “one reward” to pluralism: distributional and heterogeneous preference learning
Distributional Preference Learning (DPL).
DPL proposes to represent preferences not by a single scalar score per output but by a distribution over scores, thereby preserving uncertainty and heterogeneity that BTL-style averaging collapses Siththaranjan et al. (2024). Operationally, DPL introduces estimators that can detect when hidden context materially affects comparisons (e.g., via explained-variance diagnostics), and it demonstrates that capturing such variation can reduce downstream vulnerabilities such as jailbreak-like behaviors that exploit “regression to the mean” Siththaranjan et al. (2024). The core conceptual contribution is that heterogeneity is not noise to be averaged away; it is part of the target specification.
Pluralistic Alignment Framework (PAL) and ideal-point models.
PAL advances pluralism from a diagnostic to an architectural principle. Rather than fitting a universal preference model, PAL uses latent-variable structure (ideal-point models and mixture modeling) to learn a shared preference embedding space together with personalized or subgroup-conditioned reward functions Chen et al. (2024). This connects alignment directly to political science and psychometrics, where ideal-point models have long been used to represent heterogeneous preferences with low-dimensional latent structure. PAL shows that this approach can be efficient (leveraging penultimate-layer representations) and can few-shot generalize to new users, which is a natural analogue of “representation learning for electorates” Chen et al. (2024).
Welfare interpretation: average-case optimality versus worst-case regret.
Heterogeneity-aware alignment raises a normative question: if preferences differ, what does it mean to align to “the people”? Shirali et al. formalize alignment with user types and analyze the limits of learning a single policy under heterogeneous rewards Shirali et al. (2025). They show conditions under which the best single policy corresponds to maximizing average reward over types—a utilitarian aggregation—but also highlight that achieving this may require information that is often missing in practice (e.g., annotator identity or type labels) Shirali et al. (2025). This tension mirrors a classical divide in social choice between utilitarian welfare aggregation and protections against minority harm Harsanyi (1955); Rawls (1971).
6.4 Direct alignment objectives as tournament solutions: DPO, multi-turn extensions, and unobserved heterogeneity
DPO as “implicit reward modeling.”
DPO shows that, under a KL-regularized RLHF objective, one can optimize the policy directly using preference pairs without explicitly fitting a reward model Rafailov et al. (2023). Social-choice-theoretically, DPO can be viewed as selecting a policy that performs well in an induced tournament of outputs, where each preference comparison supplies a directed edge. This makes DPO a natural arena for importing tournament solution concepts (Condorcet, Copeland, maximal lotteries) and their differentiable surrogates Brandt et al. (2016).
Unobserved preference heterogeneity and the need for richer feedback.
Chidambaram et al. analyze DPO under unobserved heterogeneity and argue that pairwise comparisons alone can be insufficient: when different subpopulations have systematically different preferences, binary labels can confound signal in ways that cannot be disentangled without additional structure, motivating ternary or otherwise enriched feedback Chidambaram et al. (2024). This aligns with a broader identification theme: social choice outcomes depend not only on the aggregation rule but on what the ballot elicits. In alignment, “ballot design” includes the comparison protocol, annotator instructions, and the set of candidate outputs shown.
Mixture models and EM-style direct alignment.
A complementary approach is to explicitly posit latent preference types and learn a mixture of policies or a conditioned policy—conceptually moving from a single-winner election to a multi-winner or personalized system. This agenda is prominent in recent DPO extensions that use EM-style procedures to jointly infer latent types and optimize type-specific policies, providing a bridge between pluralistic alignment and direct policy optimization Chidambaram et al. (2024); Shirali et al. (2025). Such methods can be interpreted as learning a committee of aligned behaviors, with a subsequent choice rule (e.g., user-conditional selection) deciding which behavior to deploy.
Multi-turn and segment-level alignment as structured aggregation.
In interactive systems, preferences apply to trajectories, not isolated responses. SDPO proposes segment-level credit assignment for multi-turn dialogues, focusing optimization on key segments rather than individual turns (too local) or whole sessions (too coarse) Kong et al. (2025). This is naturally understood as a social choice design problem over structured objects: the “candidates” are dialogue trajectories, and the aggregator must respect temporal structure and long-horizon coherence.
6.5 Axiomatic alignment: making the aggregation principle explicit
From accidental Borda to designed aggregation rules.
The key lesson of the Borda equivalence is not that Borda is “bad,” but that alignment already implements a voting rule—often unintentionally Siththaranjan et al. (2024). The natural response is axiomatic: specify which properties we want the aggregation to satisfy and design losses or training pipelines that implement them Conitzer et al. (2024). In the alignment context, desiderata commonly include:
  • Majority responsiveness (respecting strong majority preferences),
  • Minority protection (avoiding systematic erasure),
  • Monotonicity and consistency under additional feedback,
  • Robustness to manipulation and distribution shift,
  • Transparency of what trade-off is being optimized.
Dai and Fleisig caution that translating axioms requires care: RLHF is an online, non-stationary learning system with endogenous candidate generation, so classical axioms may need adapted definitions Dai & Fleisig (2024).
Differential Voting and loss-design as rule specification.
Differential Voting provides a constructive methodology: define the aggregation rule via a differentiable loss that approximates a desired voting criterion (e.g., soft Copeland or soft Kemeny), enabling gradient-based optimization toward Condorcet-like behavior without discrete search An et al. (2026). Although introduced in the voting-rule setting, the same principle applies to alignment: if preference learning and direct alignment objectives are tournament aggregators, then loss design is mechanism design. This viewpoint also connects to the rapidly expanding ecosystem of “direct alignment” objectives (e.g., KTO and related human-aware losses), which reinterpret alignment losses through behavioral utility models Ethayarajh et al. (2024).
Constitutional AI as principle aggregation.
Constitutional AI replaces many low-level preference labels about harmfulness with a smaller set of high-level principles (a “constitution”), using model-generated critiques and revisions guided by these principles Bai et al. (2022). Social-choice-theoretically, this changes the ballot: from pairwise comparisons to normative constraints. It can be interpreted as moving from preference aggregation to judgment aggregation over principles and their instantiations, where consistency and completeness become central concerns List & Pettit (2002). The open challenge is that a constitution is itself a social choice object: whose principles, which conflicts, and what tie-breakers?
6.6 Strategic feedback, poisoning, and mechanism design for alignment pipelines
Annotators are strategic agents once they understand the aggregation rule.
A distinctive aspect of alignment is that the “electorate” (annotators, red-teamers, users) may have incentives to steer the model. DPL formalizes this by showing that under hidden context and implicit Borda aggregation, annotators can have incentives to misreport preferences to influence the learned reward function Siththaranjan et al. (2024). This reframes data collection as mechanism design: the platform must choose incentives, sampling strategies, and aggregation rules that remain robust when participants attempt to manipulate outcomes.
Content moderation and strategic adaptation as adjacent alignment mechanisms.
Strategic responses are not limited to alignment datasets. Content moderation systems induce an explicit game between a filter and users who adapt content to pass the filter. Ahmadi et al. analyze this as a mechanism design problem balancing freedom of speech against social distortion (manipulation cost), showing NP-hardness of optimal trade-offs and providing approximation and generalization guarantees from finite data Ahmadi et al. (2025). The connection to alignment is methodological: both are instances of learning under strategic adaptation, where the deployed system changes the distribution of future inputs and feedback.
Robust aggregation and pluralistic defenses.
Robustness can be pursued by (i) adopting aggregation rules that are less manipulable than Borda-like averaging (e.g., Condorcet-consistent or minimax-regret approaches), (ii) explicitly modeling preference heterogeneity so that attackers cannot exploit the “mean” behavior, and (iii) designing feedback protocols that elicit richer information (ternary or contextual labels) to reduce identification failures under heterogeneity Chidambaram et al. (2024); Siththaranjan et al. (2024). These strategies parallel robust statistics: defend the estimator (aggregation) by changing both the objective and the data-collection process.
6.7 Open Problems
OP25: From empirical audits to certified axiomatic guarantees.
Can we provide post-training certificates that an alignment procedure implements a chosen aggregation principle (e.g., Condorcet-consistency analogues) under realistic assumptions about candidate generation and distribution shift?
OP26: Minority protections without personalization collapse.
How can we design pluralistic alignment methods that protect minority values without devolving into unstable over-personalization or enabling targeted manipulation of subgroup models?
OP27: Ballot design for alignment.
What feedback primitives (pairwise, ternary, contextual rationales, principle-based critiques) are information-theoretically sufficient to identify a socially desirable aggregation outcome under heterogeneous preferences?
OP28: Strategic resistance in data collection and deployment.
How should incentives, sampling policies, and aggregation rules be co-designed so that users and annotators cannot strategically steer models toward harmful or partisan behaviors?
OP29: Constitution selection as a meta-social-choice problem.
If constitutional alignment depends on a set of principles, how should those principles be chosen, updated, and legitimately governed when societies disagree?
OP30: Bridging welfare objectives and safety constraints.
Can we reconcile welfare-theoretic aggregation (utilitarian averages, max-min fairness, regret minimization) with hard safety constraints in a way that is transparent about trade-offs and robust to misspecification?
7 Inverse Mechanism Learning: Reconstructing the Game
7.1 Section overview: differentiable forensics
In the preceding chapters, mechanisms are designed: we choose auction rules, voting rules, or allocation policies, and optimize them (often end-to-end) for revenue, welfare, or axiomatic properties. In many real systems, however, the mechanism is not directly observable. We see strategic behavior—bids in an ad exchange, routing choices in a city under tolls, platform engagement under a recommender, feature manipulation under a classifier—yet the incentive map that generated these behaviors is opaque, proprietary, or dynamically changing. This motivates an inverse problem: recover the mechanism from behavioral traces. The attached plan frames this as the “forensic tool” of differentiable social choice: rather than only building learned mechanisms, we must also audit and reconstruct the mechanisms already governing digital life.
Inverse mechanism learning differs from inverse reinforcement learning (IRL) in what is inferred. IRL aims to recover an agent’s reward function given a known environment and dynamics Ng & Russell (2000). Here, the target is the game form or incentive mechanism itself: a mapping from joint actions to (possibly individualized) payoffs, potentially unstructured and high-dimensional. This distinction matters because the mechanism can encode platform policy (e.g., an auction format), institutional constraints (e.g., subsidy schedules), or algorithmic interventions (e.g., ranking and moderation rules). It is also the hardest regime computationally: the inverse learner must reason through equilibrium selection and/or the learning dynamics of the agents. In the differentiable paradigm, this becomes “differentiate through adaptation”—unroll best-response or learning updates, backpropagate through the trajectory, and optimize a likelihood of observed behavior under candidate mechanisms. The plan highlights DIML (Differentiable Inverse Mechanism Learning) as a cutting-edge instantiation of this idea An & Du (2026).
7.2 Problem formulations: from structural econometrics to differentiable inverse games
Classical identification: auctions and structural econometrics.
The most mature predecessor of inverse mechanism learning is the structural econometrics of auctions, which asks: given observed bids (and institutional details), what can be inferred about private values, risk aversion, information, and the revenue consequences of alternative rules? Seminal results show identification of latent value distributions under first-price auctions in independent private value settings Guerre et al. (2000), and more broadly, identification across standard auction formats (first-price, second-price, English, Dutch) under flexible information structures Athey & Haile (2002). Surveys emphasize that the inverse problem is fundamentally an identification question: what features of the mechanism are uniquely recoverable from the data without heroic assumptions? Gentry et al. (2018)
However, classical auction econometrics typically assumes that the rule (e.g., first-price) is known and uses equilibrium reasoning to infer preferences or information. Inverse mechanism learning inverts the emphasis: the mechanism itself may be unknown, nonstandard, or learned by the platform. This pushes beyond parametric mechanism classes toward function approximation for payoff maps, while preserving identification discipline from econometrics.
Inverse game theory and payoff recovery.
A second predecessor is inverse game theory: infer utilities (or game parameters) that rationalize observed play, often under equilibrium assumptions. Kuleshov and coauthors provide early formalizations for succinct games, characterizing valid utilities via linear inequality systems analogous to IRL constraints Kuleshov & Schrijvers (2015). More recent work studies payoff recovery under entropy-regularized equilibria and Markov game settings, emphasizing that regularization can improve identifiability by smoothing best-responses Liao et al. (2025). Inverse mechanism learning generalizes these by allowing the “utility” to be generated by an unknown mechanism, potentially neural and non-separable across players.
Learning games to predict outcomes: structured prediction with equilibria.
A third line connects to differentiable game modeling in machine learning. Garg and Jaakkola propose parameterized games as structured predictors, mapping contextual inputs to player utilities and predicting deliberative outcomes via equilibria Garg & Jaakkola (2020). This “forward” modeling of games from data supplies tools for the inverse direction: if equilibria can be embedded in learning loops (via implicit differentiation or unrolling), then mechanism parameters can be inferred by maximizing the likelihood of observed outcomes.
7.3 Differentiating through adaptation: equilibria, learning dynamics, and unrolling
Equilibrium-based inverse inference and implicit differentiation.
One approach treats observed behavior as an equilibrium (Nash, correlated, quantal response) of an unknown mechanism. If the equilibrium map is well-behaved, one may apply the implicit function theorem to differentiate the equilibrium with respect to mechanism parameters. This is a standard move in differentiable optimization layers and is increasingly used in game-theoretic learning. Related work on differentiable games clarifies the solution concepts that gradient-based dynamics tend to find (e.g., stable fixed points vs. local Nash equilibria), which matters when “equilibrium” is selected by learning rather than by axiomatic choice Letcher et al. (2019). In Stackelberg and bilevel settings, implicit differentiation provides principled gradients through best-response maps and learning dynamics Fiez et al. (2020).
Trajectory-based inference via unrolled learning.
A second approach is closer to contemporary deep learning practice: assume agents update strategies using differentiable learning rules (e.g., no-regret updates, logit response / quantal response, policy-gradient-style learners), unroll these updates for $T$ steps, and fit mechanism parameters so that the induced action distributions match observed trajectories. This replaces a brittle equilibrium assumption with a behavioral process assumption. It also aligns with a fundamental empirical reality: many markets and platforms are not at equilibrium; they are perpetually learning systems.
A growing theory of learning in games supports this view. Unified stochastic approximation frameworks characterize long-run behavior of broad classes of learning dynamics (multiplicative weights, gradient methods, logit learning), clarifying what kinds of limit sets or stationary distributions might be expected under noise and partial information Mertikopoulos et al. (2024). For inverse problems, these results provide a “physics” model: we explain behavior by specifying how agents adapt.
7.4 DIML: likelihood-based differentiable inverse mechanism learning
The plan identifies DIML as the first general framework explicitly targeting inverse mechanism learning from multi-agent learning trajectories An & Du (2026). We summarize DIML’s core ideas because they crystallize the methodological template for this chapter.
Model class and objective.
DIML assumes an unknown incentive mechanism parameterized by $\theta$, which may itself be an unstructured neural map
$M_{\theta}:\;\mathbf{a}\mapsto(u_{1},\dots,u_{n}),$
mapping joint actions $\mathbf{a}$ to per-agent payoffs. Agents update actions (or mixed strategies) using a differentiable learning rule $\mathcal{L}$, producing a stochastic process over actions:
$\mathbf{a}_{t+1}\sim\mathcal{L}\big(\mathbf{a}_{t},\;M_{\theta},\;\xi_{t}\big).$
Given observed trajectories $\{\mathbf{a}_{t}\}_{t=0}^{T}$, DIML defines a differentiable likelihood $p_{\theta}(\mathbf{a}_{0:T})$ by unrolling the induced action distributions and maximizing $\log p_{\theta}$ over $\theta$ An & Du (2026). This yields a computationally explicit objective for reconstructing mechanisms from behavior: choose the mechanism under which the observed adaptation path is most probable.
Counterfactual payoffs and deviation reasoning.
A technical challenge is that many learning rules require comparing realized payoffs with counterfactual payoffs for deviations. DIML uses the candidate mechanism to generate these counterfactual payoffs (e.g., “what payoff would agent $i$ have obtained under action $a^{\prime}_{i}$ given $\mathbf{a}_{-i}$?”) and uses them to define gradients that align the mechanism with observed action changes An & Du (2026). This resembles regret-based training in forward mechanism design, but in reverse: rather than minimizing deviation incentives to induce truth-telling, DIML uses deviation-based signals to infer the incentives that must have been present.
Identifiability and consistency.
DIML explicitly frames which aspects of a mechanism are identifiable from trajectories. Under conditional logit response models, it establishes identifiability of payoff differences and proves statistical consistency of maximum likelihood estimation under regularity conditions An & Du (2026). This is an important conceptual contribution: in many settings, absolute payoff levels are not recoverable (only differences matter for choice), and multiple mechanisms can be behaviorally equivalent given limited observation.
Empirical domains: congestion tolling and public goods.
DIML demonstrates reconstruction on simulated environments including congestion tolling and public goods games, scaling to large anonymous games with many agents An & Du (2026). These domains are emblematic for “differentiable social choice” forensics: tolls and subsidies are mechanisms intended to align selfish behavior with social objectives, and auditing them requires inferring the effective incentives from observed behavior rather than relying on nominal policy descriptions.
7.5 Strategic opacity and fundamental limits: online strategic classification as an inverse problem
Inverse mechanism learning becomes substantially harder when agents can manipulate what we observe. Strategic classification formalizes this: a classifier (the mechanism) labels individuals based on observable features, but individuals can strategically manipulate features along a manipulation graph to obtain favorable outcomes. Ahmadi, Blum, and Yang provide fundamental bounds in the online setting, showing that strategic response changes learnability in qualitative ways and that mistake and regret bounds necessarily scale with the manipulation graph degree $\Delta$ Ahmadi et al. (2023).
While this work does not directly reconstruct a hidden mechanism, it offers inverse-learning “guardrails” in precisely the regime relevant for auditing: if observed behavior is endogenously shaped by strategic manipulation, then any inference procedure must respect lower bounds and may require stronger feedback (e.g., observing pre-manipulation states, knowing the manipulation graph, or using randomized timing models). Their separation between randomized models—whether agents respond to a distribution over classifiers or to the realized draw—highlights a broader message for inverse mechanism learning: the observational protocol is part of the mechanism. Ahmadi et al. (2023)
7.6 Applications: auditing auctions, platforms, and policy interventions
Auction forensics and format detection.
A central auditing task is detecting whether a platform is effectively running a first-price, second-price, or hybrid auction, possibly with reserve prices, bidder-specific adjustments, and dynamic ranking. Classical identification results show what can be inferred when the format is known Guerre et al. (2000); Athey & Haile (2002). Inverse mechanism learning aims to go further by inferring the effective payoff map or payment rule directly from bidding dynamics. A promising hybrid direction is to combine parametric structure (e.g., known allocation rule family) with differentiable inverse inference for unknown components (e.g., reserve schedules), enabling both identifiability (from econometrics) and flexibility (from deep learning).
Networked externalities: tolls, routing, and congestion games.
Congestion tolling is a prototypical mechanism with network constraints and learning agents. Observed route choices reflect not only the toll schedule but also bounded rationality and learning dynamics. By unrolling agent learning, inverse methods can infer the tolling mechanism that best explains adaptation, and then simulate counterfactual policies (e.g., “what would happen under marginal cost pricing?”) An & Du (2026). This connects inverse mechanism learning to inverse game models in transportation and to modern differentiable optimization for network flows.
Platform governance: recommender incentives and moderation mechanisms.
Many digital systems induce strategic user adaptation: content creators and consumers respond to ranking and moderation incentives, often trying to “game” the system. While forward mechanism design asks how to set incentives to reduce harm, inverse mechanism learning asks what incentives the platform actually implemented, potentially detecting drift, unintended bias, or deliberate policy changes. Strategic classification theory indicates that when users can manipulate features, inference and control must be robust to adversarial adaptation Ahmadi et al. (2023). This suggests a role for inverse learning as an oversight primitive: regulators or auditors could estimate the effective mechanism and evaluate it against normative criteria.
7.7 Open Problems
OP31: Scalable identifiability in high-dimensional mechanisms.
What structural assumptions (symmetry, sparsity, monotonicity, local interactions) are sufficient to make neural mechanisms identifiable from realistic behavioral traces at platform scale?
OP32: Inference under non-equilibrium, non-stationary learning.
How can inverse mechanisms be reconstructed when agents are learning with changing step sizes, shifting populations, and non-stationary objectives rather than converging to a stable equilibrium?
OP33: Joint inference of mechanism and observation protocol.
When strategic agents manipulate what is observed, how can we simultaneously infer the hidden mechanism and the manipulation/measurement process that mediates observations?
OP34: Robustness to strategic obfuscation and collusion.
How can inverse mechanism learning remain reliable when agents deliberately alter behavior to hide the mechanism (or to spoof a different one), including coordinated collusion?
OP35: Counterfactual validity and causal identification.
Under what conditions do counterfactual simulations under an inferred mechanism correspond to real-world interventions, rather than artifacts of misspecified learning dynamics?
OP36: Auditable guarantees and “forensic certificates”.
Can inverse mechanism learning produce verifiable certificates of reconstructed properties (e.g., “equivalent to first-price with reserve in this family”) that are meaningful for governance and regulation?
8 Conclusion
8.1 From axiomatic artifacts to learned institutions
This Review has traced a coherent paradigm shift: from axiomatic democracy, where mechanisms are derived as static artifacts of logic, to differentiable social choice, where mechanisms are implemented as dynamic, parameterized functions optimized from data Dütting et al. (2017; 2019); Anil & Bao (2021); Fairstein et al. (2024); Kesari & Kesari (2025); An & Du (2026). The practical motivation is clear. Classical social choice and mechanism design have produced profound conceptual tools and impossibility results—Arrow, Gibbard–Satterthwaite, and Myerson remain defining pillars Arrow (2020); Gibbard (1973); Satterthwaite (1975); Myerson (1981). Yet in many modern deployments (multi-item auctions, constrained public spending, decentralized edge aggregation, and preference-based training of frontier models), the relevant decision spaces are high-dimensional, non-convex, and distribution-dependent. In these regimes, the classical program of closed-form derivations and universal guarantees can be simultaneously too rigid (unable to exploit structure in real data) and too pessimistic (worst-case failures that may be vanishingly rare or, conversely, dangerously frequent) Hart et al. (2013); Caiata et al. (2025).
Differentiable social choice does not replace social choice theory; it re-uses it as a specification language and an audit vocabulary. In differentiable economics, incentive compatibility reappears as regret losses, Lagrangian penalties, and (in some constructions) architectural constraints derived from implementability structure Dütting et al. (2019); Ivanov et al. (2022); Hertrich et al. (2023). In neural social choice, anonymity and neutrality become permutation-equivariant inductive biases, while classical axioms become differentiable surrogates and post-hoc audits Matone et al. (2024); Hornischer & Terzopoulou (2024). In participatory budgeting, proportionality and efficiency are operationalized as loss terms (or solver-layer constraints) over knapsack-feasible portfolios, with the Method of Equal Shares (MES) serving as a principled baseline and benchmark for learned rules Peters et al. (2021); Fairstein et al. (2024); Baychkov et al. (2025). In liquid democracy, delegation graphs become architectures for decentralization and trust-weighted aggregation, with direct relevance to federated learning and edge systems Kesari & Kesari (2025); Alouf-Heffetz et al. (2025); Berinsky et al. (2023). In AI alignment, the strongest conceptual warning is that optimization is never value-neutral: standard preference-learning objectives can implicitly instantiate specific voting rules—notably Borda count—whether or not designers intend it Siththaranjan et al. (2024); Dai & Fleisig (2024). Finally, inverse mechanism learning closes the loop by turning differentiable tools toward forensics: reconstructing opaque incentive maps from observed strategic behavior to enable auditing and accountability An & Du (2026); Ahmadi et al. (2023).
8.2 Cross-cutting methodological lessons
Across domains, several methodological patterns recur.
Inductive bias is the new axiom.
Permutation symmetry (anonymity/neutrality), set structure (variable electorate size), graph structure (delegation, bipartite elections), and combinatorial feasibility (budgets, allocations) are too central to be learned reliably from scratch. Modern differentiable designs increasingly encode these properties directly via attention and graph architectures Ivanov et al. (2022); Zhang et al. (2025); Li et al. (2026), canonical embeddings from social choice Matone et al. (2024), and feasibility-enforcing layers or decompositions for constrained allocation Zhang et al. (2026); Fairstein et al. (2024).
Soft constraints are useful but brittle.
Regret minimization (approximate DSIC), monotonicity penalties, and proportionality proxies enable gradient-based optimization, but they can fail silently: a mechanism may score well on in-distribution training tests yet violate the intended principle under stronger attacks or distribution shift Hornischer & Terzopoulou (2024); You et al. (2026). This motivates a stronger research emphasis on verification, certification, and worst-case evaluation suites, not only average-case metrics.
Robustness is normative and distributional.
A recurring theme is that the most important failures often arise under misspecification: shifted bidder populations, new preference cultures, adversarial manipulation, and changing platform dynamics. The economics literature has begun to formalize these concerns via refined notions of distributional robustness that avoid “performance cliffs” when the ambiguity set is slightly wrong Ball & Kattwinkel (2024). In social choice, the analogous move is from binary axiom satisfaction to empirical violation rates under realistic distributions, which can invert classical intuitions about which rule is “more democratic” in practice Caiata et al. (2025). In security-focused work, robustness becomes adversarial training against manipulative coalitions or strategic learners Li et al. (2026); Nedelec et al. (2019).
Loss design is mechanism design.
Differential Voting crystalizes a general principle: the loss function can be treated as the explicit specification of an aggregation rule, enabling targeted approximation of hard criteria (e.g., Condorcet-style objectives) through smooth surrogates An et al. (2026). The alignment literature makes the stakes immediate: preference-learning losses can encode unintended aggregation rules, making “axiomatic alignment” a concrete engineering task rather than a philosophical aspiration Siththaranjan et al. (2024); Rafailov et al. (2023).
Auditability must be designed, not appended.
As learned mechanisms become more expressive (Transformers, GNNs), interpretability and auditability become central system requirements. Work on axiomatic audits of neural voting rules, adversarial robustness for elections, and certification of strategyproof auction networks points toward a future in which deployment requires both performance and formal or semi-formal guarantees Hornischer & Terzopoulou (2024); Li et al. (2026); Curry et al. (2020).
8.3 Future outlook: toward natural language and multi-modal mechanism design
A notable frontier is that the interface of collective decision-making is changing. Participatory budgeting and alignment already operate on unstructured text: comments, rationales, deliberation transcripts, and preference explanations. Initial studies show that LLMs can help infer preferences and produce allocations, but their feasibility reasoning and optimality are prompt- and scaffold-dependent Damle & Faltings (2025). This suggests a hybrid architecture for “natural language mechanism design”: LLMs for elicitation, summarization, and explanation, paired with certified optimization or audited rule layers for the final decision Damle & Faltings (2025); Amos & Kolter (2017). The same pattern appears in alignment: constitutional and principle-based approaches can be interpreted as changing the ballot from pairwise comparisons to normative constraints, raising meta-social-choice questions about how principles are selected and updated Bai et al. (2022); List & Pettit (2002).
At the same time, inverse mechanism learning hints at a governance toolkit for the algorithmic public sphere. If platforms can learn mechanisms by gradient descent, regulators and civil society may need methods to infer and audit those mechanisms from behavioral traces, while accounting for strategic manipulation and observation bias An & Du (2026); Ahmadi et al. (2023). In this sense, differentiable social choice is not only a design agenda; it is also an accountability agenda.